### v4

# KEN KANGDA WREN1

Abstract
Existing approaches to asset-pricing under model-uncertainty adapt classical utility-maximisation frameworks and seek theoretical comprehensiveness. We move toward practice by considering binary model-uncertainties and by switching attention from ’preference’ to ’constraints’. This decomposes economic asset-pricing into the viable pricing of model-risk and non-model risk separately such that the former has a unique and intuitive formula with convenient properties. Its parameter, dynamically conserved under model-risk inference, allows an integrated representation of ex-ante risk-pricing and bias, with ex-post price-effects that can be disentangled, through well-known anomalies, Momentum and Low-Risk, whose risk-reward curves acquire a new significance: peak-rewards measure ex-ante risk-pricing, and peak-locations, bias.
JEL Classification: C11, C12, C22, D81, G02, G12
Conflict-of-Interest Disclosure Statement
Author K.K.Wren
I have nothing to disclose.
Introduction
Motivation and Results
We study the pricing of model-risks that may be subject to uncertainty in the sense of Knight (1921) and Bewley (2002), where ’uncertainty’ refers to randomness governed by probability laws whose parameters are unknown. The aim is to move closer to practice and to examine relevant issues missed by existing literature. For our purpose, ’risk’ shall refer to randomness generally, with laws known or unknown, and ’model’, some chosen probability law.
Our approach exploits potential risk and data hierarchies, exposes model-risk as a dominant source of value gained/lost in the market, and furnishes ways to disentangle risk-pricing from bias, through price-anomalies such as Momentum (Jegadeesh and Titman (1993)) and Low-Risk (Ang et al. (2006)), without ad hoc criteria for what part of an observed drift is ’by design’ and what, ’by mistake’. The bias identified this way happens to be Status Quo Bias, a well-known behavioural trait (Samuelson and Zeckhauser (1988), Kahneman et al. (1991)), with a prominent role in Knightian Decision Theory (Bewley (2002)), making our framework a tool for the exploration of these concepts in practical and concrete terms.
Existing works adapting classical theories to incorporate model-risk come under ambiguity-aversion or robust-control, where agents perform recursive optimisation, against goals such as max-min utility, optimal smooth-ambiguity or variational preference, over the set of model choices (e.g. Ju and Miao (2012) or Machina and Siniscalchi (2014)). Related is parameter learning (e.g. Guidolin and Timmermann (2007), Collin-Dufresne et al. (2016)), carried out usually under a definitive model-risk law, without uncertainty.
Existing focus is the nature of preference/utility under model-risk and model-risk discount in a general equilibrium. Our focus is how model-risk pricing may be achieved or characterised ’minimally’, with fewest assumptions and parameters. A natural start is thus the Risk-Neutral Equivalent (RNE) formulation, given the First Theorem of Asset Pricing (FTAP) [^1] Early max-min utility struggled with FTAP-compliance (e.g. Epstein and Wang (1995)). Its modern form (e.g. Chen and Epstein (2002)) and other approaches have no such issues (e.g. Ju and Miao (2012)), especially when reformulated under the banner of smooth-ambiguity (Proposition 3.6 of Burzoni et al. (2021))..
However, viability, No Arbitrage in the FTAP sense, is too weak for most pricing tasks: there are far more viable processes than economically valid ones. Moreover, learning about model-risk makes RNE probabilities path-dependent, and so intractable, even in simple cases (proved formally here, but already observed in Guidolin and Timmermann (2007)).
One way to address such obstacles is to add incremental conditions to the FTAP baseline till RNE formulations that do provide economic asset-pricing can be characterised. It inverts the routine of ’theory making followed by viability checking’, and switches attention from ’agent preference’ to ’structural constraints’. It culminates in the decomposition of viable economic asset-pricing into separately viable components and the identification of a unique and intuitive RNE model-risk pricing formula, whose sole parameter, a constant of motion of the model-risk inferential dynamics, allows an integrated treatment of ex-ante risk-premia and bias such that their ex-post price-effects, only jointly observable, can be disentangled.
Scope and Premise
We take the existence of representative agents/beliefs for granted, and use the language of rational (Bayesian) inference to mean no more than adherence to the laws of probabilities; likewise, ’rational’ and ’expectation’ signify the same without implying ’objective correctness’, unlike in classical Rational Expectation (RE) frameworks. Three types of probabilities are thus involved: 1) objective ones governing events; 2) subjective ones representing inferential beliefs about the objective world; 3) RNE ones for pricing.
Let our model-risk $B$, of micro or macro economic origin, be binary, with outcome $b$ or $\overline{b}\not=b$, and our $B$-sure models, elementary in the sense of being model-risk free. Such $B$-sure processes may represent the familiar, which by classical RE principles must be well-modelled already. In our context therefore, only change, breaks from routines, may create model-risk (e.g. new CEO, R&D results, or regulatory/governmental/geopolitical events). Many real-world asset valuation challenges fall into this category. Generalisation to model-risks that are non-binary is outlined towards the end, by involving composite $B$-sure processes, which contain discounts for other model-risks.
Remark 1.
Some model-risks of this type may be hedged or diversified away. Our results apply only to the economically priced: any deemed worthy of being assigned at each moment a risk-premium through some recursive optimisation scheme taking said risk as an input.
Our model-risk shall be considered ’one-off’. How it ’flares up’ and the pricing of ’flare-risk’ remains implicit; that is, we study episodic ’model crises’, one at a time. This contrasts with the workhorse Markov regime-switching setting of the literature, where the pricing of the entire process is treated. Further, despite possible uncertainty and indefinite priors, our model-risk is resolvable via regular inference, albeit not before any finite horizon; this is also the case in Hansen-Sargent’s robust-control setting (e.g. Hansen et al. (2006)). Correspondingly, let horizons be sufficiently long for the meaningful development of model-risk inference, and let asset-pricing be time-homogeneous (’identical risk-levels identically priced’) during such developments (i.e. while the model-risk remains active and relevant).
The law of our model-risk, the unconditional distribution of $B$-outcomes, may be unknown. If the true unconditional probability, denoted $\mathbf{p}^{B}_{0}$ throughout, is known, model-detection (learning) and asset-pricing remain classical, with familiar conclusions. If unknown, the same may involve non-classical cognitive and economic elements. Nonetheless, asset-pricing may be formulated in terms of classical inference, with non-classical aspects reflected through preference and discount mechanisms; this is the usual, and our, strategy.
Lastly, data about model-risks (e.g. weekly/monthly indicators of economic regimes) tend to be more frequent and impactful than data about ’routine’ risks (e.g. biannual dividend). Such structures, absent in existing studies but relevant to practice, are treated explicitly here.
The setup as outlined above makes model-risk a prominent feature of asset-pricing, a major source of excess value, and so a focus of market effort and activities.
Existing Theories in Our Context
Most model-risk pricing schemes, ambiguity-aversion, robust-control and some parameter-learning, achieve model-risk discounts by in effect pessimistically distorting a reference belief (about model-risk), known as ’reference measure’, ’second-order prior’, or ’benchmark’.
The set of models for the state space $\mathcal{S}$ whose model is sought may be written as $\mathcal{I}_{1}(\mathcal{S})$. It is often identified with the space or a subspace of probability measures on $\mathcal{S}$. We see it simply as an index set, labelling model candidates. Potential laws governing the model-risk may be likewise collected into some set $\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S}))$. In our case: $\mathcal{I}_{1}(\mathcal{S})=\{b,\overline{b}\}\cong\{Q_{b},Q_{\overline{b}}\}$, one $B$-value for one model, some probability law $Q_{B}$ of state space $\mathcal{S}$, and $\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S}))\cong(0,1)$.
A reference model-risk belief thus may be identified with a member of $\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S}))$, generating learning about model-risk given the models in $\mathcal{I}_{1}(\mathcal{S})$. It is in our case some $\pi^{B}_{0}\in(0,1)$, leading to a model-risk inference process $\{\pi^{B}_{t}\}$ based on competing models $\{Q_{b},Q_{\overline{b}}\}$.
Remark 2.
Call the set $\mathcal{L}_{\mathcal{S}}:=\{\pi^{I}_{0}Q_{I}|\pi^{I}_{0}\in\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S})),\text{ }I\in\mathcal{I}_{1}(\mathcal{S})\}$ of reference beliefs about the law governing total state space $\mathcal{I}_{1}(\mathcal{S})\times\mathcal{S}$ the set of model-uncertainty, for later reference.
The max-min utility approach offers the simplest strategy: choose at any time the model with the worst outlook from the set $\mathcal{I}_{1}(\mathcal{S})$ of alternatives (in our setting it means pricing for the pessimistic $B$-outcome till $B$-risk resolves). It is a limit of the graded optimisations adopted by the other approaches, where pessimism is tempered by reference beliefs (e.g. Hansen and Sargent (2011) penalises distorted beliefs by their relative entropy to the reference).
Key to all is then the reference belief, by which model-risk pricing is defined and computed. In RE and some parameter-learning settings reference beliefs are known to be true, leading to classical risk-discounts. Under model-uncertainty, even if the reference beliefs happen to be true objectively, perceived uncertainty still triggers extra discounts. All worked examples in the uncertainty literature presume the reference beliefs used by their agents to be unknowingly objectively true; this in our context is unjustifiable.
At the other extreme, some parameter-learning stances see realised parameter-value as deterministic, meaning $\mathbf{p}^{B}_{0}\equiv 1\text{ or }0$ if adopted here; this would exclude many types of model-risks of practical interest. In our setting then, the norm is $\pi^{B}_{0}\not=\mathbf{p}^{B}_{0}\in(0,1)$.
Any assertion of equality would mask a basic issue facing all attempts to free classical RE frameworks from any built-in coincidence of beliefs and true laws: pricing theory outputs are ex-ante but what can be observed is ex-post, so tests and calibrations are impossible without additional theories of bias. Manifestations of this difficulty include ’observational equivalence’ (’is the large ex-post risk-premium due to bias after all?’; see e.g. Cecchetti et al. (2000)) and ’joint-hypothesis’ (’is an anomaly a sign of bias in the market or in one’s pricing theory?’; see e.g. Fama (2014)). We will illustrate how this issue resolves itself in our approach.
Finally, existing theories require as inputs the total risk as a whole, model and non-model risks together, aggregated across all the assets in a given market. This makes them unwieldy for bottom-up or cross-sectional enquiries, especially when the model-risk(s) can be micro, macro or of a variety of causes. Our method, with its isolation of model-risk pricing and RNE formulation, may prove more flexible and convenient in this context. It also reveals that unless a model-risk is ’consistent’, no interpretable economic model-risk pricing is possible.
Content and Organisation
Section 1 sets out the usual asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential implication, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
It transpires that the RNE formulations of asset-pricing under model-risk are constrained by informational properties (Definition 1, Lemma 1 and Remark 14) and so must follow a canonical form (Proposition 1), with familiar properties (Corollary 1) and price dynamics (Section 2.2.2). One consequence is a potential link between Status Quo Bias and well-known market-anomalies, whose characteristics separate prospective (ex-ante) market risk-pricing from bias (if any) under model-uncertainty, revealing them separately (Section 3.5.1&3.5.2).
1 Setup
1.1 Usual-Case Assets and Information Structure
For clarity and to avoid clutter, consider assets with a single payoff, some bullet-payoff $Y\in\mathbb{R}$ at horizon (or maturity) $T$, with $1\ll T\leq\infty$; let value and pricing be in logarithm, and time- or risk-free discount be exogenous and nil (with a suitable choice of numéraire). Further:
  1. 1.
    Any incremental change to bullet-payoff $Y$, if at all, is public, such that the cumulative firmed-up (’cum-div’) component $\{Z_{t}\}$ of asset-worth, with $Z_{T}\equiv Y$ by definition, has the equation of motion below, subject to binary outcome $B\in\mathcal{B}:=\{+,-\}$, but otherwise based on standard Wiener process $\{w^{Z}_{t}\}$, of natural, usual, filtration $\{\mathcal{F}^{Z}_{t}\}$,
    $dZ_{t}(B)=r^{Z}_{B,t}dt+\sigma^{Z}_{t}dw^{Z}_{t},\text{ }t\in(0,T],$ (1)
    with $\{\mathcal{F}^{Z}_{t}\}$-adapted volatility $\{\sigma^{Z}_{t}\}$ and drift $\{r^{Z}_{B,t}\}$, on probability space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$, which is the restriction to $(0,T]$ of the usual Wiener space/measure $(\mathcal{S}^{Z},W^{Z})$ generated by $\{w^{Z}_{t}\}$, with the $B$-sure law ${W}^{Z}_{B}$ of $\{Z_{t}(B)\}$ defined via $d(W^{Z}_{-})_{t}=r^{Z}_{-,t}+\sigma^{Z}_{t}dw^{Z}_{t}$ for $B=-$, and $d(W^{Z}_{+})_{t}=r^{Z}_{+,t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$ for $B=+$; write $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$.
  2. 2.
    There may be an additional $B$-informative real-vector dataflow $\{{D}_{t}(B)\}$, of similar yet $B$-conditionally independent dynamics to $\{Z_{t}(B)\}$:
    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t},$ (2)
    all as in (1) but with ’$D$’ labelling; its independence makes it price-irrelevant without model-risk (i.e. under known $B$). Combined dataflow $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ has state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and natural filtration $\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\}$, under product $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, where $\mathbf{W}_{B}:=({W}^{Z}_{B}\times{W}^{D}_{B})$.
    Remark 3.
    Current treatments ignore the make up of data $\{\mathbf{D}_{t}\}$, as if only those relevant to model-sure pricing, $\{Z_{t}\}$, exist/matter. We consider situations with data that are exclusive to model-risk (price-irrelevant once model-sure), including cases with trivial model-sure pricing data $\{Z_{t}\}$. It tackles a frequent challenge to partitioners: a crisis market unsettled by a bout of model-risk whose economics and inferential status dominate, while non-model, routine, risks remain, independently evolving or unchanging.
  3. 3.
    Given the above, continuous (predictable) inferential testing of model-risk $B$ takes place on filtered space $(\mathcal{S}^{T}_{\mathcal{B}}:=\mathcal{B}\times\mathcal{S}^{T},\text{ }\{\mathbf{F}_{t}\},\text{ }\mathbf{F}_{T};\text{ }{\pi}^{B(\cdot)}_{0}\mathbf{W}^{T}_{B(\cdot)}(\cdot))$, where ${\pi}^{B(\omega_{B})}_{0}$ is the unconditional belief in outcome $B(\omega_{B})\in\mathcal{B}$, $\omega_{B}:=(B,\omega)\in\mathcal{S}^{T}_{\mathcal{B}}$. Under the condition of $\mathbf{W}_{+}\perp\mathbf{W}_{-}$ with $\mathbf{W}_{+}|_{t}\sim\mathbf{W}_{-}|_{t}$, $\forall t<\infty$, standard testing thus resolves $B$-outcomes at guaranteed resolution-time $T(B)=\inf\big{\{}t|\mathbf{F}_{t}\ni\{B=\cdot\}\big{\}}=\infty$.
    Remark 4.
    The sequential testing of Wiener processes differing by a fixed drift is well-known. Our setup allows time-variations via absolutely continuous clock-change, making finite resolutions ($T(B)<\infty$) possible. Such issues are largely moot under economic risk-pricing (Remark 1), with $T<T(B)\leq\infty$ (see Appendix A for more).
  4. 4.
    Such testing, optimal in the sense of Neyman-Pearson Lemma, is based on likelihood-ratios (LR) $\{L^{+/-}_{t}:=\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}\}$, or equivalently, log-LR $\{l^{+/-}_{t}:=\log L^{+/-}_{t}\}$, which form Wiener processes under broad conditions; model-risk $B$ resolves iff. $\lim_{t\to T(B)}l^{+/-}_{t}=(-1)^{\mathbf{1}_{-}(B)}\cdot\infty$ (iff. $\mathbf{W}_{+}|_{T(B)}\perp\mathbf{W}_{-}|_{T(B)}$) for some $T(B)\leq\infty$. Further, standard testing is performed under the equivalent martingale measure of one of the processes (e.g that of $B=-$); henceforth let $\mathbf{W}_{-}$ be by default this base measure, under which $r^{Z}_{-,t}=0=r^{D}_{-,t}$, $\forall t$.
  5. 5.
    The above has Bayesian representations: any $L^{+/-}_{t}\in(0,\infty)$ maps bijectively and smoothly to an a posteriori belief ${\pi}^{B}_{t}\in(0,1)$. Bayes’ Rule, in terms of odds-for, $O_{f}[{\pi}^{B}_{t}]:={\pi}^{B}_{t}/\underline{\pi}^{B}_{t}$, with ()¯$\underline{(\cdot)}=1-(\cdot)$, in effect says ’new-odds $=$ old-odds $\times$ likelihood-ratios of data’:
    $O_{f}[{\pi}^{+}_{t}]=O_{f}[{\pi}^{+}_{0}]\times\exp[{l^{+/-}_{t}}]\propto{L^{+/-}_{t}}\in(0,\infty),$ (3)
    given a priori odds $O_{f}[{\pi}^{+}_{0}]\in(0,\infty)$; so inferential odds are geometric Wiener processes.
  6. 6.
    The time-$t$, $B$-sure, expected payoff $Y^{B}_{t}:={\mathbf{E}}_{B}[Y|\mathbf{F}_{t}]$ under its $B$-sure law reads ((1)):
    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathbf{F}}_{t}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=z_{t}+y_{B}(T-t,t),$ (4)
[…]
  7. 7.
[…]
1.2 The FTAP Baseline under Model-Risk
Translated to our setup, FTAP states: market $\big{(}Y,\{S_{t}\};\text{ }\mathcal{S}^{T}_{\mathcal{B}},\{\mathbf{F}_{t}\},\{\mathbf{F}_{T}\};\text{ }\pi^{B}_{0}\mathbf{W}_{B}^{T}\big{)}$, with $\{\mathbf{F}_{t}\}$-adapted asset-prices $\{S_{t}\},\text{ }S_{T}\equiv Y$, is viable iff. there is an equivalent measure $\widehat{\pi^{B}_{0}\mathbf{W}^{T}_{B}}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, the RNE measure, such that $S_{t}=\hat{\mathbf{E}}[Y\big{|}\mathbf{F}_{t}]$, $\forall t<T$, where RNE expectation $\hat{\mathbf{E}}[\cdot{|}\cdot]$ may be written as:
$S_{t}=\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{S}^{B}_{t}}:=\hat{\pi}^{+}_{t}\hat{S}^{+}_{t}+\hat{\pi}^{-}_{t}\hat{S}^{-}_{t},$ (5)
$\hat{S}^{B}_{t}:=\hat{\mathbf{E}}_{B}[Y\big{|}\mathbf{F}_{t}]:=\hat{\mathbf{E}}\big{[}Y\big{|}B,\mathbf{F}_{t}\big{]},$ (6)
with $\Sigma_{B}{(\cdot)^{B}}{(\cdot)^{B}}$ denoting convex combinations, such as the one under RNE beliefs $\hat{\pi}^{B}_{t}$, as calculated under Bayes’ Rule (3) given $\hat{\pi}^{B}_{0}$ and RNE log-LR $\hat{l}^{+/-}_{t}:=\log\frac{d\hat{\mathbf{W}}^{T}_{+}|_{t}}{d\hat{\mathbf{W}}^{T}_{-}|_{t}}$.
This market, defined by (1-6), is complete (Martingale Representation Theorem), and so the RNE measure, unique.
Viability by no means ensures economic validity (e.g. any pricing based on RNE measures of the form $\tilde{\pi}^{B}_{0}\tilde{\mathbf{W}}^{T}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ is blind to $B$-information); indeed, any pricing that is viable to any member of the equivalence class $[{\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is viable to all in the class. That is, under FTAP-viability, ’anything goes’ (almost). What then characterises the processes (and the RNE measures) that do correspond to plausible economic theories? It turns out that having economically interpretable RNE price-of-model-risk is the key.
Consider risk-premium, the gap between ex-ante price and expectation: at any $t<T$, given reference model-risk belief $\{{\pi}^{B}_{t}\}$ ((3)),
$RP_{t}:=Y_{t}-S_{t}:=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t},$ (7)
and its RNE decomposition (by inserting ±$\pm\Sigma_{B}\hat{\pi}^{B}_{t}{{Y}^{B}_{t}}$):
$RP_{t}=\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}+B\_\widehat{RP}_{t},$ (8)
$\hat{RP}^{B}_{t}:=Y^{B}_{t}-\hat{S}^{B}_{t},$ (9)
$B\_\widehat{RP}_{t}:=RP_{t}-\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}=({\pi}^{+}_{t}-\hat{\pi}^{+}_{t}){Y}^{\Delta}(t),$ (10)
where $\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}$ accounts for $B$-sure (non-model) risks, and $B\_\widehat{RP}_{t}$, model-risk $B$. The latter has standard deviation $Y^{\Delta}(t)\sigma^{\pi}_{t}$, $\sigma^{\pi}_{t}:=(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}$, which, being free of risk-pricing choices, is an intrinsic quantity. It well-defines a RNE price-of-$B$-risk:
${k}^{\hat{\pi}}_{t}:=\frac{B\_\widehat{RP}_{t}}{Y^{\Delta}(t)\sigma^{\pi}_{t}}\equiv\frac{{\pi}^{+}_{t}-\hat{\pi}^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}.$ (11)
1.3 The Economic Decomposition of Risk-Pricing under Model-Risk
Regardless of the economic pricing scheme under model-risk, classical or otherwise, the resulting asset-price $\{S_{t}\}$, with risk-premia $\{RP_{t}\}$ and RNE formulation (6-11), admits an economic decomposition of the form (16-19), provided the existence of 1) some reference model-risk beliefs $\{{\pi}^{B}_{t}\}$ and 2) $B$-sure pricing $\{S^{B}_{t}\}$ under the same scheme when $B$-sure:
$S^{B}_{t}=\check{\mathbf{E}}_{{B}}[Y{|}{\mathbf{F}}_{t}]=z_{t}+\check{\mathbf{E}}_{B}[Y-z_{t}]=z_{t}+\check{y}_{B}(T-t,t),$ (12)
$RP^{B}_{t}:=y_{B}(T-t,t)-\check{y}_{B}(T-t,t)=:RP^{B}(t)\geq 0,$ (13)
[…]
$\check{y}_{+}(T-t,t)-\check{y}_{-}(T-t,t)=:S^{\Delta}(t)>0,\text{almost surely},$ (14)
$RP^{B}(t)=\lim_{\sigma^{\pi}_{t}\to 0}\hat{RP}^{B}(t),\text{ }\text{i.e. }S^{B}_{t}=\lim_{\sigma^{\pi}_{t}\to 0}\hat{S}^{B}_{t},$ (15)
where (14) ensures consistency with asset-property Item-7, Section 1.1, and (15), internal consistency of the pricing scheme under model-risk.
Remark 5.
Usual-case model-sure pricing is simple: static noise discounted by a deterministic drift. With model-risk $B$, the only new input to economic pricing (recursive optimisation taking drivers of the expected preference-value as inputs) is the reference model-risk belief $\{{\pi}^{B}_{t}\}$.
At any $t<T$, viable pricing must satisfy $S_{t}\in(S^{-}_{t},S^{+}_{t})$, so there always exists a well-defined pricing coefficient ${A}^{B}_{t}\in(0,1)$ such that:
$S_{t}=\Sigma_{B}{A}^{B}_{t}S^{B}_{t}=S^{-}_{t}+A^{+}_{t}S^{\Delta}(t),$ (16)
$RP_{t}:=Y_{t}-S_{t}=\Sigma_{B}{\pi}^{B}_{t}{RP^{B}(t)}+B\_RP_{t},\text{ with }$ (17)
$B\_RP_{t}:=RP_{t}-\Sigma_{B}{\pi}^{B}_{t}{RP^{B}(t)}=\Sigma_{B}\pi^{B}_{t}S^{B}_{t}-S_{t}=({\pi}^{+}_{t}-A^{+}_{t}){S}^{\Delta}(t),$ (18)
$k^{A}_{t}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}\geq 0,$ (19)
where the economic price-of-$B$-risk $\{k^{A}_{t}\}$ should be free from explicit dependence on $B$-sure pricing, the latter being based on decisions in the absence of $B$-risk.
As such, once obtained, price-of-$B$-risk $\{k^{A}_{t}\}$ may generate the coefficients $\{A^{B}_{t}\}$ via (19) for asset-pricing under $B$-risk, given consistent but otherwise arbitrary $B$-sure prices $\{{S^{\prime}}^{B}_{t}\}:=\{Y^{B}_{t}\}-\{RP^{\prime B}(t)\}$, thus separating the pricing of model and non-model risks.
This is attractive if for a useful range of risk-premia, the prices so constructed:
$\{S^{\prime}_{t}\}\equiv\{S^{(RP^{\prime}|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}{S^{\prime}}^{B}_{t}\}=\{S^{(0|A)}_{t}\}-\text{ }\{\Sigma_{B}A^{B}_{t}{RP^{\prime}}^{B}(t)\},\text{ with }$ (20)
$\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}=\{Y_{t}\}-\text{ }\{k^{A}_{t}\sigma^{\pi}_{t}\cdot{Y}^{\Delta}(t)\},$ (21)
are viable; both are one way or another the sum of a martingale and a drift. Observe that all applied economic asset-pricing theories, continuous and regular near risk-neutrality, are already viable written in above form, at least for sufficiently small risk-discounts.
Remark 6.
In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, can be dialled towards near neutrality (linearity) while secondary utility/variational-control continues to exert uncertainty discounts. In classical settings, including parameter-learning, where a single utility must be optimised, the same may be achieved by replacing the payoff variable with its up-to-date model-sure expectation perturbed by a small enough noise.
The real question is then the range of viability for processes (20-21), and most relevant to us, if they have convenient RNE formulations. For this, we must characterise the economic price-of-$B$-risk $\{k^{A}_{t}\}$ ((19)) more fully. Its origin suggests the following.
Definition 1.
Asset-pricing under model-risk $B$ is said to be normal if its price-of-model-risk $\{k^{A}_{t}\}$ at any $t\ll T$: 1) depends deterministically on reference model-risk belief $\{\pi^{B}_{t}\}$ in a time-homogenous manner [^2] Economic risk-pricing, solution to some optimisation-based pricing scheme, is measurable to the natural filtration of its input, ($(\{Z_{t}\},\{\pi^{B}_{t}(\mathbf{D}_{t})\})$, and so a function of its level to date (Proposition 4.9 of Breiman (1992)). Any explicit dependence on time $t$ or $Z_{t}$ itself, if at all, is presumed weak or non-existent, at least for $t\ll T$. such that the dependence is twice-differentiable and non-negative; 2) satisfies $\lim_{\sigma^{\pi}_{t}\to 0}k^{A}_{t}=0$; 3) generates viable pricing via (20-21) under an economically relevant range of risk-premium choices.
Remark 7.
Property-1&2 are shared by most if not all existing model-risk pricing theories. Property-3 will be shown to hold for usual-case assets under model-risk.
2 Canonical RNE Asset-Pricing under Model-Risk
2.1 Viable Normal Model-Risk Pricing and its RNE measure
Proposition 1.
For usual-case assets under model-risk $B$, with reference model-risk process $\{\pi^{B}_{t}\}$ based on some reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$ from the set of model-uncertainty, normal pricing (Definition 1) is viable iff. it has this canonical form: at any $t<T$,
$S_{t}=S^{(RP|{\Pi})}_{t}=\Sigma_{B}\Pi^{B}_{t}S^{B}_{t}=S^{(0|{\Pi})}_{t}-\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t},$ (22)
where model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$ ((21)) is given by some RNE model-risk inference process $\{\Pi^{B}_{t}\}$ based on a RNE measure of the form $\Pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Proof. See Appendix B. Briefly, any viable model-risk only pricing $\{S^{(0|A)}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}$ must have drift $\{Y^{\Delta}(t)dA^{+}_{t}\}$ offset $\{dY^{-}_{t}+A^{+}_{t}dY^{\Delta}(t)\}$ in expectation under some RNE measure, possible iff. it belongs to $\mathcal{L}_{\mathcal{S}}$, under which the two drift separately vanish in expectation, due to $\{A^{+}_{t}\}$ being a twice-differentiable function of model-risk inference $\{\pi^{+}_{t}\}$ driven partly/wholly by $\{dY^{-}_{t}\}$. The viability of (22), as a martingale plus a well-behaved drift, follows.
Canonical pricing thus allows the pricing of model and non-model risks to have separate RNE formulations that are each easy to identify and apply. Note however that the overall RNE measure of viable normal asset-pricing (22) under model-risk $B$, although guaranteed to exist, remains unidentifiable in general, even for simple cases.
Remark 8.
Note that canonical model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}$ has identical RNE and economic price-of-model-risk ((11) and (19)). This indeed exhausts the set of price processes with this property (Lemma 1 and Remark 14, Appendix A). In fact any process with non-trivial risk-pricing of the non-model risks (i.e. $\{RP^{B}_{t}\}\not=0$) must have an RNE price-of-model-risk that is not a time-homogeneous deterministic function of any reference model-risk belief based on the set $\mathcal{L}_{\mathcal{S}}$ of model-uncertainty [^3] Any viable price process whose RNE measure is in $\mathcal{L}_{\mathcal{S}}$ must have the form $\{\Sigma_{B}A^{B}_{t}S^{B}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$, so $\forall t<T$, $\lim_{\sigma^{\pi}_{t}\to 0}S_{t}=Y^{B}_{t}$ since $\lim_{\sigma^{\pi}_{t}\to 0}\sigma^{\Pi}_{t}=0$, but $\lim_{\sigma^{\pi}_{t}\to 0}{S}_{t}=S^{B}_{t}=Y^{B}_{t}-RP^{B}_{t}$ given (15); thus $\{RP^{B}_{t}\}=0$.. This echos the observation that parameter-learning makes RNE formulations path-dependent (Guidolin and Timmermann (2007)).
2.2 Properties of Canonical Model-Risk Pricing
2.2.1 The 1-Parameter Price-of-Model-Risk
Corollary 1.
Canonical model-risk premium $\{B\_RP_{t}\}$ and price-of-model-risk $\{k^{\Pi}_{t}\}$ are as follows:
$B\_RP_{t}$ $=({\pi}^{+}_{t}-{\Pi}^{+}_{t}){S}^{\Delta}(t),\text{ }t<T,$ (23)
$k^{\Pi}_{t}$ $=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t},\text{ }t<T,$ (24)
with $\sigma^{\Pi}_{t}:=({\Pi}^{+}_{t}\underline{\Pi}^{+}_{t})^{\frac{1}{2}}$, and $K:=\frac{O_{f}[{\pi}^{+}_{0}]}{O_{f}[{\Pi}^{+}_{0}]}=\frac{O_{f}[{\pi}^{+}_{t}]}{O_{f}[{\Pi}^{+}_{t}]}$, a conserved quantity of risk-inference ((3)).
Remark 9.
[…]
Remark 10.
The sign of the difference process $\{{\pi}^{+}_{t}\}-\{{\Pi}^{+}_{t}\}$ is fixed ((3)); so unless model-risk impact is consistent (Item-7, Section 1.1), model-risk premia $\{B\_RP_{t}\}$ must turn negative at times: inconsistent model-risks cannot have interpretable viable economic risk-premia.
2.2.2 Connection with Familiar Price Dynamics
[…]
$S_{t}=S^{(\check{rp}|\Pi)}_{t}=S^{(0|\Pi)}_{t}-\check{rp}(t)=Y^{-}_{t}-\check{rp}(t)+S^{\Delta}(t)\Pi^{+}_{t},$ (25)
$dS_{t}(B)=dY^{-}_{t}(B)+\check{R}_{t}dt-\Pi^{+}_{j}(B)r^{Z\Delta}_{t}dt+S^{\Delta}(t)d\Pi^{+}_{t}(B).$ (26)
Given Section 1.1 and inference dynamics (A.2-A.3), we have $\{d\Pi^{+}_{t}\}$ as below, with two sources of noise, one from $\{Z_{t}\}$, given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}$ (’signal-to-noise’ ratio), one from independent $B$-informative data $\{D_{t}\}$, modelled in parallel ((2)), which can be absorbed via $\{\sigma^{l}_{t}dw^{\prime}_{t}\}=\{\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}\}$, with $(\sigma^{l}_{t})^{2}=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$:
$d\Pi^{+}_{t}(B)=\mu^{\Pi,B}_{t}+\sigma^{\Pi,l}_{t}dw^{\prime}_{t},\text{ with},$ (27)
$\mu^{\Pi,B}_{t}:=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\Big{(}{{\mathbf{1}_{+}}}(B)-{\Pi}^{+}_{t}(B)\Big{)},$ (28)
$\sigma^{\Pi,l}_{t}:=(\sigma^{\Pi}_{t})^{2}\sigma^{l}_{t}.$ (29)
The above has a drift by design relative to reference belief $\{{\pi}^{+}_{t}\}$ (of dynamics (A.3)):
$\mu^{\Pi,\pi}_{t}:=\Sigma_{B}{\pi}^{B}_{t}{\mu^{\Pi,B}_{t}}=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot(\pi^{+}_{t}-\Pi^{+}_{t})=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot k^{\Pi}_{t}\sigma^{\pi}_{t}\geq 0;$ (30)
note ${\mu^{\pi,\pi}_{t}}\equiv 0\equiv{\mu^{\Pi,\Pi}_{t}}$, confirming the martingale property of inference and model-risk pricing. Further, as inference volatility $\sigma^{\Pi,l}_{t}$ (29) drives price-volatility ((26-27)), the price-of-model-risk (24) peaks as model-risk induced price-volatility peaks.
Pulling it together, the full asset-price dynamic, a $(\Pi^{B}_{0}\mathbf{W}^{T}_{B})$-martingale plus drift $\{\check{R}_{t}\}$, reads:
$dS_{t}(B)=$ $\big{[}\check{R}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}$ (31)
$+$ $r^{Z\Delta}_{t}dt\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{j}(B)\big{)}$ (32)
$+$ $S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\sigma^{l}_{t})^{2}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{t}(B)\big{)}+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}.$ (33)
Under trivial $\{Z_{t}\}$ (firmed-up value process), the above becomes $dS_{t}(B)\propto d\Pi^{+}_{t}(B)$, making asset-pricing a pure matter of model-risk inference, which further reduces to the standard inferential test of Peskir and Shirayev (2006) under constant $\{r^{D\Delta}_{t}\}$ (signal) and $\{\sigma^{D}_{t}\}$ (noise).
Remark 11.
Practical cases often see $\check{R}_{t}\ll r^{Z\Delta}_{t}\ll S^{\Delta}(t)$: the $B$-sure dynamic (31) dominated by model-drift (32) dominated by model-inference (33); dispersion among such assets is driven by model-risk inference. It often holds also that $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$, so $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$: prices dynamics are driven by data other than those on tangible (firmed-up) values.
Averaged under model-risk belief $\{{\pi}^{+}_{t}\}$, dominant price-dynamic (33) has a classical form:
$\sigma_{t}\bigg{(}\frac{\mu_{t}}{\sigma_{t}}+dw^{\prime}_{t}\bigg{)},$ (34)
with volatility $\sigma_{t}:=\sigma^{\Pi,l}_{t}S^{\Delta}(t)$, drift $\mu_{t}:=\mu^{\Pi,\pi}_{t}S^{\Delta}(t)$, and price-of-diffusion-risk $\mu_{t}/\sigma_{t}$ ((24)):
$\frac{\mu_{t}}{\sigma_{t}}=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t}\sigma^{\pi}_{t}\sigma^{l}_{t}=\frac{K^{\frac{1}{2}}-K^{-\frac{1}{2}}}{S^{\Delta}(t)}\bigg{(}\frac{\sigma^{\pi}_{t}}{\sigma^{\Pi}_{t}}\bigg{)}\sigma_{t}.$ (35)
Expansion in $(K-1)$, given Remark 9, makes the classical relationship more explicit:
$\frac{\mu_{t}}{\sigma_{t}}\approx\bigg{(}\frac{K-1}{S^{\Delta}(t)}\bigg{)}\sigma_{t}\text{ and }\mu_{t}\approx\bigg{(}\frac{K-1}{S^{\Delta}(t)}\bigg{)}\sigma_{t}^{2}.$ (36)
Remark 12.
The proportionality factor characterises ex-ante pricing through an intuitive and direct metric of asset-pricing (Remark 9). How its level comes about is a matter of economic theory (e.g. risk-aversion). Regardless, our derivation links it firmly to ex-ante beliefs, making the above relevant to bias studies and measurements under model-uncertainty.
3 Bias and Risk-Pricing Measurement via Price Anomalies
3.1 The Joint-Hypothesis Issue
Let the $\{\mathbf{F}_{t}\}$-adapted objective conditional probabilities of model-risk outcomes be $\{\mathbf{p}^{B}_{t}\}$. The ex-post drift of our asset, $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$, assuming $B$-indifferent $B$-sure risk-premia $\{\check{rp}_{t}\}$ (Section 2.2.2) for clarity, satisfies:
$rp_{t}-\check{rp}(t)=B\_RP_{t}+(\mathbf{p}^{+}_{t}-{\pi}^{+}_{t}){S}^{\Delta}(t),\text{ }t<T.$ (37)
The LHS terms are each known/observable. The RHS terms are entangled: no conclusion on risk-pricing $\{B\_RP_{t}\}$ can be made without one on $\{\mathbf{p}^{b}_{t}\}-\{{\pi}^{b}_{t}\}$, namely, bias. For our model-risk, any realised process under a realised $B$-outcome corresponds to one run of (31-33) as if $\mathbf{p}^{b}_{t}\equiv 1\text{ or }0$. Many ’like’ observations are needed for an estimate of (37). Yet, given ’one-off’ and heterogeneous sources of model-uncertainties, how can this be done?
3.2 Assets Facing Potential Model-Change
When model-sure, our asset is a simple drift with a static noise. Upon some trigger, it may face an objective chance $\mathbf{p}^{(\cdot)}_{0}$ of entering a new state (e.g. from $B=+$ to $B=-$). Denote any new state by ’1’, any status quo, ’0’, and status quo model-sure pricing, $\{S^{\text{{\bf{0}}}}_{t}\}$.
Change-trigger may or may not follow a known law. Our interest lies in what ensues after its arrival. It amounts to treating model-risk as a compound process and ignoring the part governing how/when model-risk arises (considered known, valued and reflected in model-sure prices). The average interval between model crises, say $\lambda^{-1}$, is presumed long enough for our one-off methodology; further, consider $t\ll\lambda^{-1}$ and $t\ll T$ henceforth.
With the economic nature $sign[1{\text{{\bf{0}}}}]:=sign[Y^{1}_{t}-Y^{\bf{0}}_{t}]$ of potential change explicit, ignoring $t$-dependence of change impact $S^{\Delta}(t)$ for $t\ll T$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$, canonical pricing reads ((25)):
$S_{t}=S^{\text{{\bf{0}}}}_{t}+sign[1{\text{{\bf{0}}}}]S^{\Delta}\cdot\Pi^{1}_{t},\text{ }t\ll T,$ (38)
where the given RNE belief process $\{\Pi^{1}_{t}\}\sim\{\pi^{1}_{t}\}$ can be characterised by risk-pricing parameter $K:=(\frac{O_{f}[\pi^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]})^{sign[1{\text{{\bf{0}}}}]}$ (Corollary 1); let the market be competitive, so $K\in(1,2)$ (Remark 9).
3.3 The Representation of Status Quo Bias
The objective conditional probability process $\{\mathbf{p}^{1}_{t}\}$ governing change-risk is determined by the objective but unknown unconditional law $\mathbf{p}^{1}_{0}$ of model-risk. Its relationship against the RNE $\{\Pi^{1}_{t}\}$ and reference belief process $\{\pi^{1}_{t}\}$ can be characterised through the following:
$\rho K^{sign[1{\text{{\bf{0}}}}]}\equiv\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]},$ (39)
$\rho:=\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\pi}^{1}_{0}]};$ (40)
for instance, if a change-risk is unpriced, then $K=1$; or if it is classical, with its law $\mathbf{p}^{1}_{0}$ known and so without model-uncertainty, then $\rho=1$.
Parameter $\rho$, born of model-uncertainty, is not part of any ex-ante economic risk-discounting. The joint-hypothesis issue takes a concrete form here: observable prices reflect the product (39), and to split it into ’intention’ $K$ and ’mistake’ $\rho$ external criteria seem essential. This, as will be seen, need not be so; indeed, a high apparent bias under model-uncertainty may not be irrational either. The case of interest for us is when change is rare, $\mathbf{p}^{1}_{0}<\frac{1}{2}$, but priced as even rarer, such that (39-40) are both sizeable ($\rho\gg 1$). This ’justified strong status quo’ serves as a clear base against which ’change’ is defined and assessed (’unjustified strong status quo’, $\mathbf{p}^{1}_{0}>\frac{1}{2}$, allows easy wins to bets on ’change’ and so cannot persist).
3.4 Unconditional Model-Risk Symmetry, Dominance and Uniformity
Consider a market of a large number of assets subject to change-risks (e.g. new CEO, tort, war, recession). Let these risks be as characterised and economically priced in this market (e.g. due to their aggregate impact on market value/volatility/dispersion [^5] Dispersion is particularly intriguing in our context. As a priced factor it has been linked to market uncertainty, the business cycle and ’structural shifts’ (e.g. Demirer and Jategaonkar (2013), Kolari et al. (2021)).).
For clarity, let the unconditional all-cause probability of positive/negative economic change be even: at any time half of all potential changes in the market are positive/negative. Unconditional unevenness is addressed in Remark 17&18.
Let ’pricing intention’ $K$ ((39)) and ’apparent bias’ $\rho$ ((40)) be indifferent to the economic nature of potential change. Further, let parameter $K$, $\rho$ and change-impact $S^{\Delta}$ be uniform across the assets. Such uniformity amounts in practice to asset dispersions being driven by differences in inferential progression rather than in other aspects (Remark 11). Similarly, let all risk-inference be based on i.i.d data, with log-LR processes of fixed and uniform volatility. Note that each individual realisation of a change-outcome is considered random and independent still; while this may be a stretch for triggers such as war, it is a sensible start.
3.5 Conditional Cross-Sectional Trading
Pair-trading sorts assets into cohorts by cross-sectional statistics for long vs short positions. It amounts in our setup to conditioning by $B$-inferential events $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, the driver of dispersion. By Bayes’ Rule (3) and definition (39), the $\{\Pi^{1}_{t}=v\}$-conditioned ex-post mean-excess $rp(\pm,v):=\pm(\mathbf{p}^{1}_{t}(v)-v)S^{\Delta}$ given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change reads:
$rp(\pm,v)=\pm(\sigma^{v})^{2}\frac{1-(\rho K^{\pm 1})^{-1}}{v+\underline{v}(\rho K^{\pm 1})^{-1}}S^{\Delta},$ (41)
where $(\sigma^{v})^{2}:=v\underline{v}$, making price-volatility the driver of excess ((29) and (34)). Note that the above is invariant under label-switching $1\leftrightarrow{\text{{\bf{0}}}}$ (i.e. $v\leftrightarrow\underline{v}$, +$\leftrightarrow$-+ ↔ - and $\rho\leftrightarrow\rho^{-1}$).
Remark 13.
With no uncertainty, $\rho=1$, the above is the ex-ante risk-premium when $\{\Pi^{1}_{t}=v\}$ ((24)): in classical markets, all the ’strategising’ merely earns what the market deems fair.
Despite excess (41) being driven by volatility, conditioning by $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$ under strong status quo amounts to ranking by momentum (past performance): $\{\Pi^{1}_{t}=v\}$ means $\{\Pi^{1}_{t}\}$ going from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v$ and so an excess-to-date of ±$\pm(v-\Pi^{1}_{0})S^{\Delta}$ ((38)). The associated model-uncertainty ($\propto(\sigma^{v})^{2}$) is a priori low and then evolves with dataflow. The window of opportunity for pair-trading is given by (C.8), Appendix C.2.
Drifts are hard to measure (e.g. Merton (1980)). Volatility trading is a good practical alternative. Volatility events, by (26-29), take the form $\{\Pi^{1}_{t}=v\}\cup\{\Pi^{1}_{t}=\underline{v}\}$, $v\in(0,\frac{1}{2}]$, and the $\{\sigma^{\Pi}_{t}=v\underline{v}\}$-conditioned ex-post mean-excess $rp_{t}^{\sigma}(v)$ averages over the $sign[1{\text{{\bf{0}}}}]$ of potential change and over the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-events in the volatility cohort:
$rp_{t}^{\sigma}(v):=\frac{\sum_{\pm}\mathbb{P}_{t}(\pm,v)rp(\pm,v)+\sum_{\pm}\mathbb{P}_{t}(\pm,\underline{v})rp(\pm,\underline{v})}{\mathbb{P}_{t}(v)+\mathbb{P}_{t}(\underline{v})},$ (42)
where $\mathbb{P}_{t}(\cdot)=\sum_{\pm}\mathbb{P}_{t}(\pm,\cdot):=\sum_{sign[1{\text{{\bf{0}}}}]}\mathbb{P}_{t}\big{(}\{sign[1{\text{{\bf{0}}}}]=\cdot\}\cap\{\Pi^{1}_{t}=\cdot\}\big{)}$.
The relevant likelihoods (ratios) are determined by the underlying log-LR dynamics, whose cross-sections are Normal, making closed expressions possible. As a by-product, the ratio functions predict an association of volatility with ’negative momentum’ (Remark 17), a well-known phenomenon (e.g. Ang et al. (2006) and Wang and Xu (2015)).
The risk-reward curves of momentum trading (41) and volatility trading (42) are both concave in $v$ and vanishing as $v\underline{v}\to 0$.
3.5.1 Momentum Effect and Status Quo Bias
Conditional excess (41) exhibits Momentum, a key price-anomaly (Jegadeesh and Titman (1993)): if $\rho K^{\pm 1}\gg 1$, excess-to-date ±$\pm(v-\Pi^{1}_{0})S^{\Delta}$, due to RNE beliefs $\{\Pi^{1}_{t}\}$ rising from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v>\Pi^{1}_{0}$, persists on average in the same direction. The effect applies to $\rho K^{\pm 1}\ll 1$ also: past excess due to RNE beliefs $\{\Pi^{1}_{t}\}$ falling from $\Pi^{1}_{0}\approx 1$ to $v<\Pi^{1}_{0}$ means an average future excess of the same sign; indeed, label-switching here defines a new parameter $\rho^{-1}>1$. Momentum thus vanishes only under $\rho\approx 1$ (little model-uncertainty).
The effect grows with $\rho$, which by definition (40) measures bias against change. From (41), momentum trading earns its best realisable excess $rp^{\pm mo}_{max}$ at $v^{\pm mo}_{max}$:
$v^{\pm mo}_{max}:=\frac{1}{(\rho K^{\pm 1})^{\frac{1}{2}}+1},\text{ }rp^{\pm mo}_{max}:=\pm\frac{(\rho K^{\pm})^{\frac{1}{2}}-1}{(\rho K^{\pm})^{\frac{1}{2}}+1}S^{\Delta}.$ (43)
Peak-profitability reads $\frac{1}{2}(rp^{+mo}_{max}-rp^{-mo}_{max})$, by pairing cohorts of conditioning-momentum $v^{\pm mo}_{max}$. As a function of bias $\rho$, risk-reward curve (41) and its peak (43) imply rising rewards from rising momentum up to a point, beyond which profitability falls, as seen in Ang et al. (2006).
3.5.2 Low-Risk Effect and Status Quo Bias
Low-Risk Effect (e.g. Ang et al. (2006)) turns the classical dictum ’high risk high reward’ on its head. Momentum risk-reward (41) already hints at it: peak-reward (43) is in the low-volatility region if bias $\rho$ is sizeable. The effect is clearer under volatility-conditioning (42), whose peak-location and -size can be found by routine methods (Appendix C). The solution at $\rho=1$ (no bias) and leading-order solutions at $\rho\gg 1$ (high bias) are revealing and may be given by the same expression; the peak-location $v^{\sigma}_{max}$ and -size $rp^{\sigma}_{max}$ are:
$v^{\sigma}_{max}:=\frac{1}{\rho+1},\text{ }rp^{\sigma}_{max}:=\frac{1}{2}\frac{K-1}{K+1}S^{\Delta}.$ (44)
At $\rho=1$ the above is exact, by which $v^{\sigma}_{max}=\frac{1}{2}$, where price-volatility peaks, thus confirming ’high risk high reward’ in the absence of model-uncertainty or bias.
There is otherwise a low-risk effect: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, corresponding to low risk if $\rho\gg 1$; peak volatility, $v=\frac{1}{2}$, is rewarded poorly ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$. Note the separation of risk-pricing $K$ and bias $\rho$: the latter is revealed by peak-location $v^{\sigma}_{max}$, while the former, by peak-reward $rp^{\sigma}_{max}$, which is that set ex-ante for peak-risk (Remark 9). No external rules are needed to tell ’intended pricing’ from ’unintended bias’.
4 Summary and Discussion
We investigate the risk-neutral equivalent formulation of model-risk pricing. Explicit price dynamics under binary model-risks with potentially unknown laws are derived, exposing model-risk as a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk. It offers an integrated and practicable approach to these highly relevant and interconnected topics.
We make use of a feature of standard hypothesis testing (Lemma 1) that ’informationally equivalent’ tests must be essentially identical. This makes risk-neutral equivalent formulations under model-risk path-dependent and intractable. On the other hand, it limits viable economic asset-pricing (Remark 1) under model-risk to just one form (Proposition 1), leading to an intuitive risk-pricing formula with intuitive properties (Corollary 1).
The sole risk-pricing parameter is associated with a conservation law of inference and is a product of ’intention’ (preference) and ’mistake’ (difference between objective and subjective laws). Ex-ante (’intended’) model-risk pricing, classical or with ambiguity-aversion/robust-control, can be captured by our parameter $K\in[1,\infty)$ (39), with $K=1$ corresponding to ’no model-risk discounts’. Its derivation reveals that it sets the intended gain-loss ratio of the asset at peak model-risk, suggesting $K\in[1,2)$ in competitive markets (Remark 9).
The difference between true model-risk law and the reference model-risk belief on which asset-pricing is based can be captured by our parameter $\rho\in(0,\infty)$ (40), with $\rho=1$ corresponding to ’no difference’ (the classical setting), and $\rho\not=1$, ’bias’. It gives expression to Status Quo Bias and allows its effects to be disentangled from risk-pricing (Section 3.5.1-3.5.2).
The above is interesting also in light of the inertia axiom of Bewley (2002), which says, in effect, that to update an existing decision by incorporating a new state subject to Knightian uncertainty, the status quo is kept unless an alternative that is better in all scenarios exists. For our model-risk, it suggests inertial pricing, one that is piece-wise free of model-risk pricing, jumping between alternative models when appropriate, despite data that can resolve model-risk via regular inference. Asset-prices under max-min utility already have this characteristic, albeit with a built-in pessimism absent in inertial pricing. Both are viable in our setup, as models are given by probability laws that are equivalent before risk-resolution. The theoretical and practical aspects of this topic will be examined in a follow-up study.
Our approach may be extended to model-risks characterisable as $(N+1)$-tuples of independent binary outcomes ${B}^{N+1}:=(B_{0}B_{1}...B_{N})$, $N\in\mathbb{N}$, through iteration. It may be applied first to the $2^{N}$ pairs of ${B}^{N+1}$-sure, elementary, price processes, then to the resulting $2^{N-i}$ pairs of ${B}^{N+1-i}$-sure composites, $i=1,...,N$, each discounting the $i$-tuple model-risk $(B_{0}B_{1}...B_{i-1})$.
APPENDIX
A Properties of Inferential Hypothesis Testing
  1. 1.
    Standard Continuous-Time Inference. Any continuous process with independent increments are essentially Wiener. By definition (Item-4, Section 1.1), the log-LR process under i.i.d data, in terms of some constant $\sigma^{l}>0$ (the ’signal-to-noise’ of data), reads:
    $dl^{+/-}_{\tau}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l})^{2}}{2}d\tau+\sigma^{l}dw_{\tau},$ (A.1)
    with $\{w_{\tau}\}$ denoting a standard Wiener process. Time-variation may be modelled via an absolutely continuous clock-change $t\mapsto\tau(t)$. The image process $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$, again Wiener and with absolutely continuous time-varying parameters (e.g. Kallsen (2006)), along with inferential beliefs, an Ito process of it ((3)), read as follows:
    $dl^{+/-}_{t}(B)=\mu^{l^{+/-}}_{t}(B)dt+\sigma^{l}_{t}dw_{t},\text{ }\text{ }\mu^{l^{+/-}}_{t}(B):=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l}_{t})^{2}}{2};$ (A.2)
    $\frac{d\pi^{+}_{t}(B)}{(\sigma^{\pi}_{t})^{2}}=\big{(}\mathbf{1}_{+}(B)-\pi^{+}_{t}(B)\big{)}(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}.$ (A.3)
  2. 2.
    Regular Inference. Its defining feature is the predictable resolution of risky outcomes. In usual situations (e.g. i.i.d data), risk-resolution occurs at infinity. Clock-change allows continuous-time regular inference to have finite resolution-times, by compactifying uniform regular inference to bring resolution $T_{\tau}(B)=\infty$ forward to some $T(B):=\tau^{-1}(\infty)<\infty$. The resulting log-LR process diverges as $t\to T(B)$, but the associated beliefs and so asset-pricing remain finite and continuous under Bayes’ Rule (3).
  3. 3.
    Complete Markets and Unhedgeable Model-Risks. Market completeness excludes unpredictable resolutions (price-jumps) and so demands regular inference. In this setting, to be unhedgeable and thus priced economically (Remark 1), model-risk $B$ cannot be resolvable before horizon $T$ (so $T(B)>T$); this is the default case for model-risk studies. The mixed case of probabilistic resolution, ($\mathbb{P}\big{(}\{T(B)\leq t+s\}|F_{t}\big{)}\in(0,1)$, $\forall t<T$, $\forall s\in(0,T-t]$, which may occur in continuous-time complete markets so long as $T(B)$ is predictable, involves treating the unhedgeable part of the model-risk only.
  4. 4.
    Non-Resolution and Novikov’s Condition. For any log-LR process based on equivalent Wiener measures on the Wiener space $\mathcal{S}$ of total data, written as $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ in the financial context, one being some RNE version of the other, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d\hat{\mathbf{W}}_{B}|_{t}(\cdot)}{d\mathbf{W}_{B}|_{t}(\cdot)}<\infty$ almost surely, and hence a finite total variance ((A.2)). That is, the non-resolution of $B$-outcomes (Item-4, Section 1.1) equates to (A.2-A.3) satisfying Novikov’s Condition, and vice versa.
  5. 5.
    Adjacency. Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a constant as $t\to T(B)$, since:
    $\hat{l}^{+/-}_{t}-l^{+/-}_{t}=\log{\frac{d\hat{\mathbf{W}}_{+}|_{t}}{d\mathbf{W}_{+}|_{t}}}-\log{\frac{d\hat{\mathbf{W}}_{-}|_{t}}{d\mathbf{W}_{-|_{t}}}},\text{ }\forall t\in\mathbb{R}^{+},$ (A.4)
    where the RHS is convergent under equivalence.
    Lemma 1.
    No adjacent log-LR processes, unless identical up to a constant difference, can be informationally equivalent, that is, be deterministically and time-homogenously related to one another, nor can any inferential processes derived from them.
    Proof. It is a feature of inference (and due to Bayes’ Rule). It can be verified concretely via Ito’s Lemma here, given (A.2) and restrictions to twice-differentiable maps between processes (as in Proposition 1). Any such mapping $g:{l}^{+/-}_{t}\in\mathbb{R}\mapsto\hat{l}^{+/-}_{t}=g(l^{+/-}_{t})\in\mathbb{R}$, $t\in(0,T(B))$, must solve the ordinary differential equation:
    $g^{\prime\prime}=-(-1)^{\mathbf{1}_{+}(B)}(g^{\prime}-1)g^{\prime}.$ (A.5)
    If $B=+$, then $g^{\prime}e^{-g}=g^{\prime}(0)e^{-g(0)}e^{-\mathbf{Id}}$ and so $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; divergence $\lim_{t\to T(+)}\hat{l}^{+/-}_{t}=\infty$ requires $g^{\prime}(0)=1$; If $B=-$, then $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$ and $\lim_{t\to T(-)}\hat{l}^{+/-}_{t}=-\infty$ imply the same.
    Remark 14.
    Given identical LR up to a constant factor, inference process $\{\hat{\pi}^{+}_{t}\}$, a $(\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B})$-martingale by definition, is also a $(\hat{\pi}^{B}_{0}\mathbf{W}_{B})$-martingale ((3)), meaning $\hat{\mathbf{W}}_{B}=\mathbf{W}_{B}$ (Martingale Representation Theorem and so completeness of the Wiener basis): thus, only tests based on identical underlying beliefs can be informationally equivalent.
B PROOF OF PROPOSITION 1
Proof. Part I. Viable model-risk only pricing $\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}_{t}\}$, given twice-differentiable coefficients $\{A^{+}_{t}(\pi^{+}_{t})\}$, must have the canonical form.
Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d(\mathbf{W}^{T}_{-})_{t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{Z}_{t}\}$ being independent standard Wiener processes (Item-2&4, Section 1.1), we have:
$dS^{(0|A)}_{t}=\big{[}($ $\mathbf{1}_{+}-A^{+}_{t})r^{Z\Delta}_{t}dt\text{ }+\text{ }\sigma^{Z}_{t}dw^{Z}_{t}\big{]}\text{ }+\text{ }$ (B.1)
$(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\partial_{\pi}A^{+}_{t}\cdot\bigg{[}($ $\mathbf{1}_{+}-\pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+(\sigma^{\pi}_{t})^{2}\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}(\sigma^{l}_{t})^{2}dt\text{ }+\text{ }\sigma^{l}_{t}dw^{\prime}_{t}\bigg{]},$ (B.2)
with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’ ratio) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), and $\sigma^{lD}_{t}$ likewise, from $B$-conditionally independent data $\{D_{t}(B)\}$, and so $\sigma^{l}_{t}dw^{\prime}_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
Without risk-pricing, $\{A^{+}_{t}\}=\{\pi^{+}_{t}\}$, we have trivial viability under reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}$:
$dS^{(0|\pi)}_{t}=\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})r^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}.$ (B.3)
With viable model-risk only pricing, under some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$: we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$ and the following relationship, by (8-11) and (19),
$\pi^{+}_{t}-A^{+}_{t}=k^{{A}}_{t}\sigma^{\pi}_{t}=k^{\hat{\pi}}_{t}\sigma^{\pi}_{t}+\frac{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{RP}^{B}_{t}}{{Y}^{\Delta}(t)}=(\pi^{+}_{t}-\hat{\pi}^{+}_{t})+\frac{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{RP}^{B}_{t}}{{Y}^{\Delta}(t)}.$ (B.4)
Remark 15.
Reference model-risk beliefs $\{{\pi}^{B}_{t}\}$ and any RNE $\{\hat{\pi}^{B}_{t}\}$ are ’adjacent’ (Item-5, Appendix A), implying that at any fixed $t<T$ for $\sigma^{\pi}_{t}\ll\frac{1}{2}$, by (23-24) when $\{{\hat{\pi}}^{B}_{t}\}=\{\Pi^{B}_{t}\}$, we have $\mathcal{O}({\pi}^{B}_{t}-{\hat{\pi}}^{B}_{t})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and so $\mathcal{O}({k}^{\hat{\pi}}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$ for any RNE price-of-model-risk ((11)). In turn, by Lemma 1, to balance the LHS and RHS of (B.4), we have: $\mathcal{O}({\hat{RP}^{B}_{t}})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}({k}^{A}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$, as well as $\mathcal{O}(\partial_{\pi}A^{B}_{t})=\mathcal{O}(\partial^{2}_{\pi}A^{B}_{t})=\mathcal{O}(1)$ and $\mathcal{O}(d({\hat{RP}^{B}_{t})})=\mathcal{O}(d(\sigma^{\pi}_{t})^{2})=\mathcal{O}((\sigma^{\pi}_{t})^{4})$.
Consider first the case of trivial $\{D_{t}\}$. The drift of $\{dS^{(0|A)}_{t}\}$ is nil in expectation under some RNE measure $\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T}$, where $d(\hat{W}^{Z}_{B})_{t}=\hat{r}_{B,\text{ }t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$, $\hat{r}_{B,\text{ }t}dt=-d(\hat{RP}^{B}_{t})$, $t<T$.
Being the RNE measure, the $(\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T})$-expectation of (B.1) must offset that of (B.2). Write that of (B.1) as $\nu_{t}:=(\hat{\pi}^{+}_{t}-A^{+}_{t})r^{Z\Delta}_{t}-\Sigma_{B}\hat{\pi}^{B}_{t}\hat{r}_{B,\text{ }t}$ and note $\hat{\pi}^{+}_{t}-\pi^{+}_{t}=(\hat{\pi}^{+}_{t}-A^{+}_{t})-(\pi^{+}_{t}-A^{+}_{t})$, viability thus requires:
$\frac{-\nu_{t}}{r^{Z\Delta}_{t}}=(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\partial_{\pi}A^{+}_{t}\cdot\frac{(\sigma^{lZ}_{t})^{2}}{r^{Z\Delta}_{t}}\bigg{[}(\sigma^{\pi}_{t})^{2}\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}-({\pi}^{+}_{t}-A^{+}_{t})-\frac{-\nu_{t}}{r^{Z\Delta}_{t}}\bigg{]}.$ (B.5)
By Remark 15, we have $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$ and $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift is possible only if each side of the above vanishes on its own, meaning:
$\frac{\partial_{\pi}^{2}A^{+}_{t}}{2\partial_{\pi}A^{+}_{t}}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\sigma^{\pi}_{t})^{2}}.$ (B.6)
This has solutions of the form $\{A^{+}_{t}\}=\{\Pi^{+}_{t}\}$ only, where $\{\Pi^{+}_{t}\}$ is any model-risk inference based on any measure $\Pi^{B}_{0}W_{B}^{Z}|_{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$. The same can be obtained from the LHS of (B.5).
Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d(\hat{\mathbf{W}}_{B}^{T})_{t}=\hat{r}_{B,\text{ }t}dt+d(\mathbf{W}_{-}^{T})_{t}$, $\hat{{r}}_{B,\text{ }t}dt=-d(\hat{{RP}}^{B}_{t})$. To conclude, the RNE measure $\{S^{(0|A)}_{t}\}$ must have the canonical form: $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Part II. Viability of canonical normal pricing $\{S^{(RP|{\Pi})}_{t}\}=\{{S}^{(0|{\Pi})}_{t}\}-\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$.
Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}(RP^{B}(t))$ that stems from $B$-sure risk-pricing:
$dS^{(RP|\Pi)}_{t}=\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\check{r}^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}+R^{(\cdot)}_{t}dt,$ (B.7)
where $\check{r}^{Z\Delta}_{t}dt:=-{d}{}S^{\Delta}(t)$, with $S^{\Delta}(t)\equiv Y^{\Delta}(t)-RP^{+}(t)+RP^{-}(t)$. As such, it is viable so long as this extra drift $\{R^{B}_{t}dt\}$ can be ’absorbed’ into the known $(\Pi^{B}_{0}\mathbf{W}_{B}^{T})$-martingale $\{d{S}^{(0|{\Pi})}_{t}\}$, that is, whenever
$\frac{{R^{\pm}_{t}}/{S^{\Delta}(t)}}{{\sigma^{Z}_{t}}/{S^{\Delta}(t)}+{(\sigma^{\Pi}_{t})^{2}{\sigma^{lD}_{t}}\sqrt{1+\text{F}^{Z/D}_{t}}}}$
is square-integrable up to horizon $T$ (Novikov’s Condition), where, for convenience, we have written $(\sigma^{l}_{t})^{2}=(\sigma^{lD}_{t})^{2}(1+\text{F}^{Z/D}_{t})$, with $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}$.
C THE MOMENTUM AND LOW-RISK FORMULAE
For the assets of Section 3.2-3.4, all the distributions required stem from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
$\mathcal{N}\bigg{(}\mu^{l}_{[t]}(B)=\frac{(-1)^{B+1}}{2}(\sigma^{l})^{2}t,\text{ }(\sigma^{l}_{[t]})^{2}=(\sigma^{l})^{2}t\bigg{)},\text{ }B\in\{0,1\}.$ (C.1)
The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices ((38)) then derive via change-of-variable:
$l=H_{\Pi\pm}+\log\circ\text{ }O_{f}[v],$ (C.2)
$H_{\Pi\pm}:=\log\circ\text{ }O_{f}[\underline{\Pi}^{1}_{0\pm}]=H_{\mathbf{p}}+\log(\rho K^{\pm 1}),$ (C.3)
$H_{\mathbf{p}}:=\log\circ\text{ }O_{f}[\underline{\mathbf{p}}^{1}_{0}],$ (C.4)
for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, with $dl=(v\underline{v})^{-1}dv$ ((3)) and $O_{f}[\Pi^{1}_{0\pm}]=O_{f}[\mathbf{p}^{1}_{0}]/(\rho K^{\pm 1})$ ((39)) given $sign[1{\text{{\bf{0}}}}]=\pm$. The $H$-variables above are inferential milestones: $H_{\Pi\pm}$ is the log-LR hurdle for event $\{\Pi^{1}_{t}\geq\frac{1}{2}\}$, and $H_{\mathbf{p}}$, for $\{\mathbf{p}^{1}_{t}\geq\frac{1}{2}\}$.
C.1 Tracking Inferential Progress and Bias Dominance
The degree of $B$-certainty as data accumulate may be assessed in the usual way given Normal distributions. The objective degree is tracked by some $|C^{\mathbf{p}}_{t}|$, where, with $t_{\mathbf{p}}:=2H_{\mathbf{p}}/(\sigma^{l})^{2}$,
$C^{\mathbf{p}}_{t}(B):=\frac{H_{\mathbf{p}}-\mu^{l}_{[t]}(B)}{\sigma^{l}_{[t]}}=\frac{1}{2}\sigma^{l}_{[t]}\cdot\big{(}{(-1)^{B}}+\frac{t_{\mathbf{p}}}{t}\big{)}.$ (C.5)
Certainty is high for $\sigma^{l}_{[t]}\ll 1$ (little data) or $\sigma^{l}_{[t]}\gg 1$ (lots of data), with $|C^{\mathbf{p}}_{t}|$ bottoming at 00 if and when $H_{\mathbf{p}}=\mu^{l}_{[t]}(B)$, that is, if and when $t=(-1)^{B+1}t_{\mathbf{p}}$ has a solution.
For the price-implied RNE inference process $\{\Pi^{1}_{t}\}$, the same indicator, denoted $|C^{\Pi}_{t\pm}|$, reads:
$C^{\Pi}_{t\pm}(B):=$ $\frac{H_{\Pi\pm}-\mu^{l}_{[t]}(B)}{\sigma^{l}_{[t]}}=\frac{1}{2}\sigma^{l}_{[t]}\cdot\big{(}{(-1)^{B}}+\frac{t_{\Pi\pm}}{t}\big{)}$ (C.6)
$=$ $\frac{1}{2}\sigma^{l}_{[t]}\cdot(\frac{t_{\rho}}{t}\pm\frac{t_{K}}{t})+C^{\mathbf{p}}_{t}(B),$ (C.7)
[…]
Bias creates a ’burden of proof’ $t_{\rho}$, and $t_{\rho}/t$ tracks how it is overcome. We consider cases of positive objective hurdle $H_{\mathbf{p}}>0$ and bias domination $H_{\Pi\pm}-H_{\mathbf{p}}\gg 0$, so $t_{\mathbf{p}}>0$ and $t_{\rho}\gg t_{K}$, where ’objective burden’ $t_{\mathbf{p}}$ is presumed ’non-extreme’ for non-trivial inference/excess.
C.2 The Window of Opportunity
Excess-profit opportunities occur when data become sufficient for the conditional probabilities of change to be meaningful but insufficient for subjective inference to overcome bias:
$\big{\{}\frac{t_{\mathbf{p}}}{t}\ll 1\big{\}}\text{ }\bigcap\text{ }\big{\{}\frac{t_{\rho}}{t}\gg 1\big{\}};$ (C.8)
the first demand means data dominance over objective hurdle and the second, bias over data.
Remark 16.
Consider for example profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ or $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ ((44)), with $t$ fixed and inside the above window, and $u>0$ large enough for the second event to be outside. Their respective probability densities are $\propto(t)^{-1/2}e^{-|C^{\mathbf{p}}_{t}|^{2}/2}$ and $\propto(t\pm u)^{-1/2}e^{-|C^{\mathbf{p}}_{(t\pm u)}|^{2}/2}$. Thus, in-window events dominate, with a log-LR of $\propto\big{(}\log[1\pm u/t]+|C^{\mathbf{p}}_{(t\pm u)}|^{2}-|C^{\mathbf{p}}_{t}|^{2}\big{)}\gg 1$, which diverges as $(1-u/t)\to 0$ or $(1+u/t)\to\infty$, with $t$ fixed.
C.3 Momentum and Volatility Mixtures
The probabilities relevant to volatility-conditioned mean-excess (42) obey the rule:
$\mathbb{P}_{t}(\Pi^{1}_{t}=v)\mathbb{P}_{t}(\pm|\text{ }\Pi^{1}_{t}=v)\equiv\mathbb{P}_{t}(\pm)\mathbb{P}_{t}(\Pi^{1}_{t}=v|\text{ }\pm),$ (C.9)
with $\mathbb{P}_{t}(\Pi^{1}_{t}=v|\text{ }\pm)$ from Normal distribution (C.1) via (C.2). The likelihood-ratio of $\{\Pi^{1}_{t}=\underline{v}\}$ vs $\{\Pi^{1}_{t}=v\}$, $v\in(0,\frac{1}{2}]$, given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change, is:
$R_{t|\pm}(\underline{v}/v)(B):=\frac{\mathbb{P}_{t}(\underline{v}|\pm)}{\mathbb{P}_{t}(v|\pm)}(B)=({\underline{v}}/{v})^{-(t_{\Pi\pm}/{t})-(-1)^{B}},\text{ }B\in\{0,1\}.$ (C.10)
Note that $\{\Pi^{1}_{t}=v\}$-events dominate as long as bias does (i.e. $t_{\rho}/{t}\gg 1$) regardless; its dependence on actual $B$-outcomes is weak.
To compute (42) we also need the mix-ratio function $M_{t|v}:={\mathbb{P}_{t}(+|v)}/{\mathbb{P}_{t}(-|v)}$ of the event-set $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, between member-events with $\{sign[1{\text{{\bf{0}}}}]=+\}$ and with $\{sign[1{\text{{\bf{0}}}}]=-\}$.
By (C.9) we have $M_{t|v}={\mathbb{P}_{t}(v|+)}/{\mathbb{P}_{t}(v|-)}$, since $\mathbb{P}_{t}(\pm)=\frac{1}{2}$ by setup. Hence:
$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}\cdot K^{-(t_{\mathbf{p}}/t)-(-1)^{B}},$ (C.11)
and likewise the volatility-conditioned ratio $M_{t|v\underline{v}}:={\mathbb{P}_{t}(+|v\underline{v})}/{\mathbb{P}_{t}(-|v\underline{v})},\text{ }v\in(0,\frac{1}{2}]$:
$M_{t|v\underline{v}}=M_{t|v}\frac{1+R_{t|+}(\underline{v}/v)}{1+R_{t|-}(\underline{v}/v)}.$ (C.12)
Note the source of conditionality: risk-pricing $K$; if $K=1$, conditional mixes equal the unconditional mix; if $K\in(1,2)$, they are perturbations of the unconditional one.
Remark 17.
Both mix-ratios (C.11-C.12) are monotone declining, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$. That is, for $\rho\gg 1$, in the window (C.8) of opportunity, when trading events are most abundant, the observed mix at high volatility can be highly negative vs the unconditional level. Mix returns to its unconditional state as $t\to\infty$ when model-risk resolves. Any uneven unconditional background merely introduces a constant factor in (C.11-C.12); all above statements remain valid.
C.4 Peak-Reward Location and Size for Volatility-Conditioned Trading
Focusing first on the $\{\Pi^{1}_{t}=v\}$-contributions to volatility-conditioned mean-excess (42), given mix-function (C.11) and condition (C.8), it has a unique optimum, (44), with $\mathcal{O}(\frac{(K-1)H_{\mathbf{p}}}{(\sigma^{l})^{2}t})$-errors; the accuracy of solution (44) improves with data accumulation.
Further, given (C.10) and under condition (C.8), the $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (42) are no more than $\mathcal{O}(\rho^{-t_{\rho}/{t}})$. That is, the leading order solution (44) is not affected by these contributions while $t_{\rho}/t\gg 1$, i.e. while bias dominates.
Remark 18.
For uneven unconditional-mix, $\mathbb{P}_{t}(\pm)\not=\frac{1}{2}$, the same applies. Compared to even solution (44), its peak moves left (right), and size, down (up), for unevenness that tilts negatively (positively); peak-location is capped by $v^{+mo}_{max}$ ((43)), and peak-size, by $rp^{+mo}_{max}$ ((43)).
Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-part of (42) are invariant under $v\leftrightarrow\underline{v}$ switching ((41) and (C.10-C.12)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming ’high risk high reward’ under classical conditions.