### v6

# KEN KANGDA WREN1

Abstract
Existing approaches to asset-pricing under model-uncertainty adapt classical utility-maximisation frameworks and seek theoretical comprehensiveness. We move toward practice by considering binary model-risks and by emphasizing ’constraints’ over ’preference’. This decomposes viable economic asset-pricing into that of model and non-model risks separately, leading to a unique and convenient model-risk pricing formula. Its parameter, a dynamically conserved constant of model-risk inference, allows an integrated representation of ex-ante risk-pricing and bias such that their ex-post impacts are disentangled via well-known anomalies, Momentum and Low-Risk, whose risk-reward patterns acquire a fresh significance: peak-reward reveals ex-ante risk-premia, and peak-location, bias.
JEL Classification: C11, C12, C22, D81, G02, G12
Conflict-of-Interest Disclosure Statement
Author K.K.Wren
I have nothing to disclose.
Introduction
Motivation and Results
We study the pricing of model-risks that may be subject to uncertainty in the sense of Knight (1921) and Bewley (2002), where ’uncertainty’ refers to randomness governed by probability laws whose parameters are unknown. The aim is to move closer to practice and to examine relevant issues missed by existing literature. For our purpose, ’risk’ shall refer to randomness generally, with laws known or unknown, and ’model’, some chosen probability law.
Our approach exploits potential risk and data hierarchies, exposes model-risk as a dominant source of value gained/lost in the market, and furnishes ways to disentangle risk-pricing from bias, through price-anomalies such as Momentum (Jegadeesh and Titman (1993)) and Low-Risk (Ang et al. (2006)), without ad hoc criteria for what part of an observed drift is ’by design’ and what, ’by mistake’. The bias identified this way happens to be Status Quo Bias, a well-known behavioural trait (Samuelson and Zeckhauser (1988), Kahneman et al. (1991)), with a prominent role in Knightian Decision Theory (Bewley (2002)), making our framework a tool for the exploration of these concepts in practical and concrete terms.
Existing works adapting classical theories to incorporate model-risk come under the heading of ambiguity-aversion or robust-control, where agents perform recursive optimisation, against goals such as max-min utility, optimal smooth-ambiguity or variational preference, over the set of model choices (e.g. Ju and Miao (2012) or Machina and Siniscalchi (2014)). Related also is parameter learning (e.g. Guidolin and Timmermann (2007), Collin-Dufresne et al. (2016)), often performed under some known model-risk law.
Existing focus is the nature of preference/utility under model-risk and the resulting general equilibrium. We focus here instead on how model-risk pricing might be done or characterised ’minimally’, in terms of assumptions and parameters. A natural start is thus the Risk-Neutral Equivalent (RNE) formulation, given the First Theorem of Asset Pricing (FTAP) [^1] Early max-min utility struggled with FTAP-compliance (e.g. Epstein and Wang (1995)). Its modern form (e.g. Chen and Epstein (2002)) and other approaches have no such issues (e.g. Ju and Miao (2012)), especially when reformulated under the banner of smooth-ambiguity (Proposition 3.6 of Burzoni et al. (2021))..
The RNE approach to model-risk has been seldom discussed, for two good reasons. Firstly, viability, No Arbitrage in the FTAP sense, is too weak for most purposes: there are far more viable price processes than economically valid ones. Secondly, parameter learning makes RNE probabilities path-dependent and intractable, even in simple cases, as observed by Guidolin and Timmermann (2007). We overcome these difficulties by tracing their origins and examining the consequences of the constraints they impose on asset-pricing.
The key turns out to be whether model-risk pricing is a standard Ito process of the model-risk inference process, or equivalently, whether the economic price-of-model-risk has an identical RNE form. It leads to a unique and intuitive model-risk pricing formula, whose parameter, a constant of motion of model-risk inference, allows an integrated treatment of ex-ante risk-premia and bias such that their ex-post effects, only jointly observable, may be disentangled. The RNE approach can apply at the aggregate or the individual-asset level, thereby potentially expanding the data universe helpful to and addressable by model-risk/parameter-learning studies, which are currently closed to stock-level or cross-sectional data.
Scope and Premise
The existence of representative agents/beliefs is taken for granted, so the language of rational (Bayesian) inference is used, to mean no more than adherence to the laws of probabilities; likewise, ’rational’ and ’expectation’ signify the same without implying ’objective truthfulness’, unlike in classical Rational Expectation (RE) frameworks. We thus work with three types of probabilities: 1) the objectively true ones governing events; 2) the subjective ones representing inferential beliefs about the objective world; and 3) the RNE ones for pricing.
Let our model-risk $B$ have two outcomes, $b$ or $\overline{b}\not=b$, and our $B$-sure models, elementary (’model-risk free’). Such $B$-sure processes represent ’routines’ that by classical RE principles are well-modelled; only change creates model-risk (e.g. new CEO, R&D results, macroeconomic, regulatory or geopolitical events). Non-binary situations that can be handled within the binary framework emerge quite naturally, as will be demonstrated in an application addressing typical questions encountered in practice.
Remark 1.
Some model-risks of this type may be hedged or diversified away. Our results apply only to the economically priced: any deemed worthy of being assigned at each moment a risk-premium through some recursive optimisation scheme taking said risk as an input.
Our model-risk shall be considered one-off. How it ’flares up’ and the pricing of ’flare-risk’ remains implicit; that is, we study episodic ’model crises’, one at a time. This contrasts with the workhorse Markov regime-switching setting of the literature, where the pricing of the entire process is treated. Further, despite possible uncertainty and indefinite priors, our model-risk is resolvable via regular inference, albeit not before any finite horizon; this is also the case in Hansen-Sargent’s robust-control setting (e.g. Hansen et al. (2006)). Correspondingly, let horizons be sufficiently long for the meaningful development of model-risk inference, and let asset-pricing be time-homogeneous (’identical risk-levels identically priced’) during such developments (i.e. while the given model-risk remains active and relevant).
The law of our model-risk, the unconditional distribution of $B$-outcomes, may be unknown. If the objectively true unconditional probability, denoted $\mathbf{p}^{B}_{0}$ throughout, is known, model-detection (learning) and asset-pricing remain classical, with familiar conclusions. If unknown, the same may involve non-classical cognitive and economic elements. Nonetheless, asset-pricing may be formulated in terms of classical inference, with non-classical aspects reflected through preference and discount mechanisms; this is the usual, and our, strategy.
Lastly, data on model-risk (e.g. weekly/monthly indicators of economic regimes) tend to be more frequent and impactful than data on ’routine’ risks (e.g. biannual dividend). Such structures, absent in existing studies but relevant to practice, are treated explicitly here.
The setup as outlined above makes model-risk a prominent feature of asset-pricing, a major source of excess value, and so a focus of market effort and activities.
From the Perspectives of Standard Model-Risk Studies
Most model-risk pricing schemes, ambiguity-aversion, robust-control and some parameter-learning, achieve model-risk discounts by in effect pessimistically distorting a reference belief (about model-risk), known as ’reference measure’, ’second-order prior’, or ’benchmark’.
The set of models for the state space $\mathcal{S}$ whose model is sought may be written as $\mathcal{I}_{1}(\mathcal{S})$. It is often identified with the space or a subspace of probability measures on $\mathcal{S}$. We see it simply as an index set, labelling model candidates. Potential laws governing the model-risk may be likewise collected into some set $\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S}))$. In our case: $\mathcal{I}_{1}(\mathcal{S})=\{b,\overline{b}\}\cong\{Q_{b},Q_{\overline{b}}\}$, one $B$-value for one model, some probability law $Q_{B}$ of state space $\mathcal{S}$, and $\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S}))\cong(0,1)$.
A reference model-risk belief thus may be identified with a member of $\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S}))$, generating learning about model-risk given the models in $\mathcal{I}_{1}(\mathcal{S})$. It is in our case some $\pi^{B}_{0}\in(0,1)$, leading to a model-risk inference process $\{\pi^{B}_{t}\}$ based on competing models $\{Q_{b},Q_{\overline{b}}\}$.
Remark 2.
The set $\mathcal{L}_{\mathcal{S}}:=\{\pi^{I}_{0}Q_{I}|\pi^{I}_{0}\in\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S})),\text{ }I\in\mathcal{I}_{1}(\mathcal{S})\}$ of reference beliefs about the law of total state space $\mathcal{I}_{1}(\mathcal{S})\times\mathcal{S}$ may be called ’the set of model-uncertainty’, for convenience.
The max-min utility approach offers the simplest solution: choose at any time the model with the worst outlook from the set $\mathcal{I}_{1}(\mathcal{S})$ of alternatives (meaning in our setting choosing the pessimistic $B$-outcome till $B$-risk resolves). It is a limit of the graded optimisations used by others, where pessimism is tempered by reference beliefs (e.g. Hansen and Sargent (2011) penalises distorted beliefs by their relative entropy to a chosen benchmark).
Key to all is then the reference belief, by which model-risk pricing is defined and computed. In RE and some parameter-learning settings reference beliefs are known to be true, leading to classical risk-discounts. Under model-uncertainty, even if the reference beliefs happen to be true objectively, perceived uncertainty still triggers extra discounts. All worked examples in the ambiguity-aversion/robust-control literature presume the reference beliefs of their agents to be unknowingly objectively true; this in our context is hard to justify.
At the other extreme, some parameter-learning stances see realised parameter-value as deterministic, meaning $\mathbf{p}^{B}_{0}\equiv 1\text{ or }0$ if adopted here; this would exclude many types of model-risks of practical interest. In our setting then, the norm is $\pi^{B}_{0}\not=\mathbf{p}^{B}_{0}\in(0,1)$.
Any assertion of equality would mask a basic issue facing all attempts to free classical RE frameworks from any built-in coincidence of beliefs and true laws: pricing theory outputs are ex-ante but what can be observed is ex-post, so tests and calibrations are impossible without additional theories of bias. Manifestations of this ’impasse’ include ’observational equivalence’ (’is the high realised risk-premium observed due to bias after all?’; see e.g. Cecchetti et al. (2000)) and ’joint-hypothesis’ (’is an anomaly a sign of bias in the market or in one’s pricing theory?’; see e.g. Fama (2014)). We show how this issue may resolve itself in our approach.
Finally, existing theories require ’total risk’ as a whole as inputs, model and non-model risks together, aggregated across all relevant assets in a given market. They are unwieldy for bottom-up (individual-asset) or cross-sectional enquiries, especially when the model-risk(s) can be micro-economic or heterogeneous in origin. Our method, with its isolation of model-risk pricing and RNE formulation, is more flexible and convenient for such applications.
Content and Organisation
Section 1 sets out the usual asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential application, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
We find that the RNE formulations of asset-pricing under model-risk are constrained by economic, informational, necessities (Definition 1 and Lemma 1) and so have a canonical form (Proposition 1), resulting in familiar properties (Corollary 1) and dynamics (Section 2.2.2). One implication is a potential link between Status Quo Bias and well-known anomalies, whose characteristic risk-reward curves reveal and separate ex-ante risk-premia from bias under model-uncertainty (Section 3.4.1&3.4.2).
1 Setup
1.1 The Usual-Case Asset and Information Structure
For clarity, consider an usual-case asset: it has just one, bullet, payoff, $Y\in\mathbb{R}$, at some horizon or maturity $T$ much longer than unit-time, so $1\ll T<\infty$; let value and price be in logarithm, and time- and risk-free discount exogenous and nil (in a suitable numéraire). Further:
  1. 1.
    Any incremental change (if at all) to bullet-payoff $Y$ is such that the cumulative firmed-up (’cum-div’) asset-worth $\{Z_{t}\}$, with $Z_{T}\equiv Y$, follows the dynamics below, based on a standard Wiener process $\{w^{Z}_{t}\}$, with natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to binary outcomes $\{B=b\}$ (e.g. ’high’ or ’low’) labelled by $b\in\mathcal{B}:=\{+,-\}$,
    $dZ_{t}(B)=r^{Z}_{B,t}dt+\sigma^{Z}_{t}dw^{Z}_{t},\text{ }t\in(0,T],$ (1)
    where its $B$-sure drifts $\{r^{Z}_{B,t}\}$ and $B$-independent volatilities $\{\sigma^{Z}_{t}\}$ are both $\{\mathcal{F}^{Z}_{t}\}$-predictable. The underlying $B$-sure probability space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ is the restriction to $(0,T]$ of the Wiener space/measure $(\mathcal{S}^{Z},W^{Z}_{B})$ associated with process $\{Z_{t}(B)\}$, that is, $dW^{Z}_{B,t}=dZ_{t}(B)$. The above allows the sequential testing of Wiener processes of differing drifts [^2] Text-book sequential tests (e.g. Peskir and Shirayev (2006)) concern processes of fixed drift and volatility. Our setup allows time-variations achieved through an absolutely continuous clock-change, or equivalently, $\{\mathcal{F}^{Z}_{t}\}$-predictable drifts and volatilities. For more background see Appendix A..
  2. 2.
    The economic impact of model-risk $B$ thus derives from drift difference $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$; let it be consistent: $Y^{\Delta}(t):=\int_{t}^{T}r^{Z\Delta}_{s}ds>0$ almost surely, at any $t<T$. This is intuitive, also necessary for viable and economically valid risk-pricing, as will be seen (Remark 8).
  3. 3.
    There may be a purely $B$-informative real-vector dataflow $\{{D}_{t}(B)\}$ that is $B$-conditionally independent but otherwise similar to $\{Z_{t}(B)\}$, based on a standard Wiener process $\{w^{D}_{t}\}$,
    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t}.$ (2)
    It by definition is price-irrelevant in the absence of model-risk (when $B$-sure). Combined, total dataflow $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ has state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and natural filtration $\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\}$, under the product $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, with $\mathbf{W}_{B}:={W}^{Z}_{B}\times{W}^{D}_{B}$.
    Remark 3.
    Existing studies ignore potential structures within dataflow ($\{\mathbf{D}_{t}\}$), as if only data relevant to tangible, firmed-up value, $\{Z_{t}\}$, exist/matter. By considering data that are exclusive to model-risk (price-irrelevant once model-sure), including when tangible-value data, $\{Z_{t}\}$, are absent or trivial, we tackle a frequent challenge facing partitioners: a crisis market unsettled by a bout of model-risk whose economics dominate, while non-model, routine, risks remain, independently evolving or unchanging.
  4. 4.
[…]
  5. 5.
    Tests based on likelihood-ratios (LR) $L^{+/-}_{t}:=\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}$, or log-LR $l^{+/-}_{t}:=\log L^{+/-}_{t}$, $\forall t<T$, are optimal in the sense of Neyman-Pearson Lemma. Model-risk $B$ resolves iff. $\exists T_{\mathcal{B}}\leq\infty$ such that $\lim_{t\to T_{\mathcal{B}}}l^{+/-}_{t}=(-1)^{\mathbf{1}_{-}(B)}\cdot\infty$ (iff. $\mathbf{W}_{+}|_{T_{\mathcal{B}}}\perp\mathbf{W}_{-}|_{T_{\mathcal{B}}}$). Further, such inferential tests are performed under the equivalent martingale measure of one of the processes (e.g that of $B=$-). Henceforth let $\mathbf{W}_{-}$ be this base measure, under which $r^{Z}_{-,t}=0=r^{D}_{-,t}$, $\forall t$.
  6. 6.
    The above has a Bayesian representation: any $L^{+/-}_{t}\in(0,\infty)$ maps bijectively and smoothly to an a posteriori belief ${\pi}^{B}_{t}\in(0,1)$. Bayes’ Rule, in terms of odds-for, $O_{f}[{\pi}^{B}_{t}]:={\pi}^{B}_{t}/\underline{\pi}^{B}_{t}$, with $\underline{(\cdot)}=1-(\cdot)$, says in effect ’new-odds == old-odds $\times$ likelihood-ratios of data’:
    $O_{f}[{\pi}^{+}_{t}]=O_{f}[{\pi}^{+}_{0}]\times\exp[{l^{+/-}_{t}}]\propto{L^{+/-}_{t}}\in(0,\infty),$ (3)
    given a priori odds $O_{f}[{\pi}^{+}_{0}]\in(0,\infty)$. That is, inferential odds are exponentiated log-LR processes; the latter, $\{l^{+/-}_{t}\}$, are Wiener processes under broad conditions ((A.2-A.3)).
  7. 7.
    Given $Z_{t}=z_{t}$, $t<T$, the $B$-sure expected payoff $Y^{B}_{t}:={\mathbf{E}}_{B}[Y|\mathbf{F}_{t}]\equiv{\mathbf{E}}_{B}[Y|\mathcal{F}^{Z}_{t}]$ reads ((1-2)):
    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathcal{F}^{Z}_{t}}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=z_{t}+y_{B}(T-t,t),$ (4)
    where $y_{B}(T-t,t)-y_{B}(T-t-t^{\prime},t+t^{\prime})=\mathbf{E}_{B}[\int_{t}^{t+t^{\prime}}dZ_{s}(B)]$, $\forall t^{\prime}\leq T-t$, and so $y_{B}(T-t,t)=\int_{t}^{T}{r}^{Z}_{B,s}ds$, as by definition $\{Y^{B}_{t}\}$ is a martingale and $y_{B}(0,T)\equiv 0$. Note also that this relates to model-risk impact $Y^{\Delta}(t)$ via $Y^{\Delta}(t)\equiv Y^{+}_{t}-Y^{-}_{t}\equiv{y}_{+}(T-t,t)-{y}_{-}(T-t,t)$.
1.2 The FTAP Baseline under Model-Risk
[…]
Such FTAP-viable asset-pricing under model-risk $B$ may be written always as follows:
$S_{t}=\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{S}^{B}_{t}}:=\hat{\pi}^{+}_{t}\hat{S}^{+}_{t}+\hat{\pi}^{-}_{t}\hat{S}^{-}_{t},$ (5)
$\hat{S}^{B}_{t}:=\hat{\mathbf{E}}_{B}[Y\big{|}\mathbf{F}_{t}]:=\hat{\mathbf{E}}\big{[}Y\big{|}B,\mathbf{F}_{t}\big{]},$ (6)
where $\Sigma_{B}{(\cdot)^{B}}{(\cdot)^{B}}$ denotes averages weighted by $B$-outcomes, such as by RNE beliefs $\hat{\pi}^{B}_{t}$ under Bayes’ Rule (3) given $\hat{\pi}^{B}_{0}$ and RNE log-LR $\hat{l}^{+/-}_{t}:=\log\frac{d\hat{\mathbf{W}}^{T}_{+}|_{t}}{d\hat{\mathbf{W}}^{T}_{-}|_{t}}$.
Viability by no means ensures economic validity (e.g. any pricing based on RNE measures of the form $\tilde{\pi}^{B}_{0}\tilde{\mathbf{W}}^{T}\sim{\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ is blind to $B$-informative data); indeed, any pricing that is viable to any member of the equivalence class ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}]$ is viable to all in the class. In this sense, ’anything goes’ under FTAP-viability. What characterises the processes and the RNE measures that do correspond to plausible economic theories?
For answers, consider risk-premium, the gap between ex-ante pricing and expectation, non-negative under economic asset-pricing: at $\forall t<T$, given reference model-risk belief $\{{\pi}^{B}_{t}\}$,
$RP_{t}:=Y_{t}-S_{t}=\Sigma_{B}{\pi}^{B}_{t}Y^{B}_{t}-S_{t}\geq 0.$ (7)
It has RNE decomposition (by inserting $\pm\Sigma_{B}\hat{\pi}^{B}_{t}{{Y}^{B}_{t}}$):
$RP_{t}=\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}+B\_\widehat{RP}_{t},$ (8)
$\hat{RP}^{B}_{t}:=Y^{B}_{t}-\hat{S}^{B}_{t},$ (9)
$B\_\widehat{RP}_{t}:=RP_{t}-\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}=({\pi}^{+}_{t}-\hat{\pi}^{+}_{t}){Y}^{\Delta}(t),$ (10)
where $\Sigma_{B}\hat{\pi}^{B}_{t}{\hat{RP}^{B}_{t}}$ accounts for $B$-sure (non-model) risks, and $B\_\widehat{RP}_{t}$, model-risk $B$. The latter has standard deviation $Y^{\Delta}(t)\sigma^{\pi}_{t}$, $\sigma^{\pi}_{t}:=(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}$, which, being free of risk-pricing choices, is an intrinsic quantity. It well-defines a RNE price-of-$B$-risk:
${k}^{\hat{\pi}}_{t}:=\frac{B\_\widehat{RP}_{t}}{Y^{\Delta}(t)\sigma^{\pi}_{t}}\equiv\frac{{\pi}^{+}_{t}-\hat{\pi}^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}.$ (11)
Decomposition (8-11) exists whenever the RNE measure exists, that is, for all viable pricing schemes. However, its elements, $\{\hat{S}^{B}_{t}\}$ and $\{{k}^{\hat{\pi}}_{t}\}$, need not correspond to economically interpretable entities. It turns out that establishing such a correspondence, albeit indirectly, is the key to finding economically valid asset-prices (out of mere viable ones).
1.3 The Economic Decomposition of Risk-Pricing under Model-Risk
Just as any viable price process admits a RNE decomposition of the embedded risk-premia, any economic pricing scheme under model-risk, classical or otherwise, with asset-price $\{S_{t}\}$, risk-premia $\{RP_{t}\}$ and RNE formulation (6-11), admits an economic decomposition, (16-19). It is unique and well-defined provided the existence of 1) a reference model-risk belief process $\{{\pi}^{B}_{t}\}$; 2) viable $B$-sure pricing $\{S^{B}_{t}\}$ produced under the same scheme when $B$-sure:
$S^{B}_{t}=\check{\mathbf{E}}_{{B}}[Y{|}{\mathbf{F}}_{t}]=z_{t}+\check{\mathbf{E}}_{B}[Y-z_{t}]=z_{t}+\check{y}_{B}(T-t,t),$ (12)
$RP^{B}_{t}:=Y^{B}_{t}-S^{B}_{t}=y_{B}(T-t,t)-\check{y}_{B}(T-t,t)=:RP^{B}(t)\geq 0,$ (13)
where $\check{\mathbf{E}}_{{B}}[\cdot|\cdot]$ denotes expectations under its $B$-sure RNE measure $\check{\mathbf{W}}^{T}_{B}\sim{\mathbf{W}}^{T}_{B}$, leading to RNE outlook $\check{y}_{B}(T-t,t)=\int_{t}^{T}\check{r}_{B,s}ds$ as in (4), and $B$-sure risk-premium $RP^{B}(t)$ satisfying: $\forall t\leq T$,
$\check{y}_{+}(T-t,t)-\check{y}_{-}(T-t,t)=:S^{\Delta}(t)>0,\text{ almost surely},$ (14)
$S^{B}_{t}=\lim_{\sigma^{\pi}_{t}\to 0}\hat{S}^{B}_{t},\text{ }\text{i.e. }RP^{B}(t)=\lim_{\sigma^{\pi}_{t}\to 0}\hat{RP}^{B}(t),$ (15)
where (14) ensures consistency with asset-property Item-2, Section 1.1, while (15), the internal consistency of the pricing scheme under model-risk $B$.
Remark 4.
Model-sure usual-case pricing is simple: static noise compensated by a predictable drift. Adding model-risk, the only new input to economic pricing (recursive optimisation taking drivers of the expected preference-value as inputs) is the reference model-risk belief $\{{\pi}^{B}_{t}\}$.
[…]
$S_{t}=\Sigma_{B}{A}^{B}_{t}S^{B}_{t}=S^{-}_{t}+A^{+}_{t}S^{\Delta}(t),$ (16)
$RP_{t}:=Y_{t}-S_{t}=\Sigma_{B}{\pi}^{B}_{t}{RP^{B}(t)}+B\_RP_{t},\text{ with }$ (17)
$B\_RP_{t}:=RP_{t}-\Sigma_{B}{\pi}^{B}_{t}{RP^{B}(t)}=\Sigma_{B}\pi^{B}_{t}S^{B}_{t}-S_{t}=({\pi}^{+}_{t}-A^{+}_{t}){S}^{\Delta}(t),$ (18)
$k^{A}_{t}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\pi^{+}_{t}\underline{\pi}^{+}_{t})^{\frac{1}{2}}}\geq 0,$ (19)
which parallels RNE decomposition (8-11), but pricing coefficients $\{A^{B}_{t}\}\in(0,1)$, unlike $\{\hat{\pi}^{B}_{t}\}$, need not be a conditional-probability process. We show shortly that they in fact must be.
The economic price-of-$B$-risk $\{k^{A}_{t}\}$ should have no explicit dependencies on $B$-sure prices, the latter being decisions in the absence of $B$-risk. Thus, once devised, process $\{k^{A}_{t}\}$ can generate asset-prices under $B$-risk via (19) and (16) given any consistent but otherwise arbitrary $B$-sure pricing $\{{S}^{B}_{t}\}\equiv\{Y^{B}_{t}\}-\{RP^{B}(t)\}$ ((13)), thus separating the pricing of model and non-model risks. This is attractive if for a useful range of risk-discount choices, the asset-pricing so achieved,
$\{S^{(RP|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}{S}^{B}_{t}\}=\{S^{(0|A)}_{t}\}\text{ }-\text{ }\{\Sigma_{B}A^{B}_{t}{RP}^{B}(t)\},\text{ with }$ (20)
$\{S^{(0|A)}_{t}\}:=\{\Sigma_{B}A^{B}_{t}Y^{B}_{t}\}=\{Y^{-}_{t}\}+\{A^{+}_{t}Y^{\Delta}(t)\}=\{Y_{t}\}\text{ }-\text{ }\{k^{A}_{t}\sigma^{\pi}_{t}\cdot{Y}^{\Delta}(t)\},$ (21)
is viable. Observe that both above can be the sum of a martingale and a well-behaved drift, and that all applied economic pricing theories, continuous and regular near risk-neutrality, are already viable when written in this form at least for small enough risk-discounts.
Remark 5.
In ambiguity-aversion settings, the primary utility function, enforcing risk-aversion, and the secondary utility/variational-control, enforcing uncertainty-aversion, can be dialled independently and arbitrarily close to neutrality (linearity), to achieve the above separation. In classical and some parameter-learning settings, with a single utility, the above obtains by replacing payoffs with their model-sure expectations perturbed by suitably small noise.
The real question is then the range of viability for (20-21) and if they have tractable RNE formulations. For this, let us characterise the economic price-of-$B$-risk $\{k^{A}_{t}\}$ better:
Definition 1.
Asset-pricing under model-risk $B$ is said to be usual if its price-of-$B$-risk $k^{A}_{t}$, (19), is a time-homogenous twice-differentiable function of its model-risk belief [^3] Economic risk-pricing, as the solution to some optimisation scheme, is measurable to the natural filtration of its input $(\{Z_{t}\},\{\pi^{B}_{t}(\mathbf{D}_{t})\})$; so it is at any time $t$ a function of $(Z_{t},\pi^{B}_{t}(\mathbf{D}_{t}))$ (Proposition 4.9 of Breiman (1992)). Any explicit dependence on $t$ or $Z_{t}$, if at all, is presumed here weak or non-existent, at least for $t\ll T$. $\pi^{B}_{t}$, $\forall t\ll T$.
Remark 6.
This makes $\{k^{A}_{t}\}$ a time-homogeneous Ito process of risk-inference $\{\pi^{B}_{t}\}$, a common feature of economic risk-pricing theories: priced risks are captured as state-variables so that the resulting risk-pricing is their function and hence path-independent with respect to them. It will be seen (via Proposition 1 and Corollary 1) that this is responsible for all the ’nice’ properties of $\{k^{A}_{t}\}$: fixed-signed, with $\lim_{\sigma^{\pi}_{t}\to 0}k^{A}_{t}=0$ (ensuring (15)), and generating prices (20-21) that follow familiar dynamics and allow a wide range of viable risk-premium choices.
2 Canonical RNE Asset-Pricing under Model-Risk
2.1 Viable, Usual, Model-Risk Pricing
Proposition 1.
Under model-risk $B$ and model-risk inference $\{\pi^{B}_{t}\}$ based on a reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$ that belongs to the set $\mathcal{L}_{\mathcal{S}}$ of model-uncertainty (Remark 2), any usual pricing (Definition 1) of any usual-case asset is viable iff. it has the canonical form below: $\forall t<T$,
$S_{t}=S^{(RP|{\Pi})}_{t}=\Sigma_{B}\Pi^{B}_{t}S^{B}_{t}=S^{(0|{\Pi})}_{t}-\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t},$ (22)
where model-risk only asset-pricing $S^{(0|{\Pi})}_{t}=\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}$ (see (21)) is given by a RNE model-risk inference process $\{\Pi^{B}_{t}\}$ under a RNE measure of the form $\Pi^{B}_{0}\mathbf{W}^{T}_{B}\in\mathcal{L}_{\mathcal{S}}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Proof. Appendix B. Briefly, viable pricing $S^{(0|A)}_{t}=Y^{-}_{t}+A^{+}_{t}Y^{\Delta}(t)$ must see $dY^{-}_{t}+A^{+}_{t}dY^{\Delta}(t)$ offset $Y^{\Delta}(t)dA^{+}_{t}$ in expectation under a RNE measure, possible iff. the two expectations vanish separately due to $\{A^{+}_{t}\}$ being an Ito process of $\{\pi^{+}_{t}\}$, thus limiting the RNE to the form stated. The viability of (22) then follows from the ’good behaviour’ of $\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$
Canonical model-risk only pricing $\{S^{(0|{\Pi})}_{t}\}$, with $\{RP^{B}_{t}\}=0$, has identical economic and RNE price-of-model-risk ((19) and (11)); indeed any pricing with this property must be model-risk only (with $\{RP^{B}_{t}\}=0$) [^4] Any such pricing must have its RNE measure in $\mathcal{L}_{\mathcal{S}}$ and so have the form $\{\Sigma_{B}A^{B}_{t}S^{B}_{t}\}=\{\Sigma_{B}\Pi^{B}_{t}Y^{B}_{t}\}$. Hence, $\forall t<T$, $\lim_{\sigma^{\pi}_{t}\to 0}S_{t}=Y^{B}_{t}$ since $\lim_{\sigma^{\pi}_{t}\to 0}\sigma^{\Pi}_{t}=0$, but $\lim_{\sigma^{\pi}_{t}\to 0}{S}_{t}=S^{B}_{t}=Y^{B}_{t}-RP^{B}_{t}$ given (15); thus $\{RP^{B}_{t}\}=0$.. Given Definition 1 and Remark 6, this means that viable prices (16) that have non-trivial $B$-sure risk-pricing (with $\{RP^{B}_{t}\}\not=0$) must have a RNE price-of-model-risk that is path-dependent with respect to any model-risk process $\{\pi^{B}_{t}\}$ derived from the set $\mathcal{L}_{\mathcal{S}}$ of model-uncertainty. This formalises the observation of Guidolin and Timmermann (2007) that learning makes RNE probabilities path-dependent and so intractable.
The source of this difficulty is basically (15) (see footnote-4), given Lemma 1 and the fact that the operation of $B$-conditioning and of risk-neutralising (reference beliefs such as $\pi^{B}_{0}\mathbf{W}^{T}_{B}$) do not commute in general. Canonical pricing (22) does not remove this intrinsic issue but avoids it by separating the RNE pricing of model and non-model risks, each by itself easy and tractable. The total RNE measure of (22), though guaranteed, remains intractable.
2.2 Properties of Canonical Model-Risk Pricing
2.2.1 The 1-Parameter Price-of-Model-Risk Formula
Corollary 1.
Canonical price-of-model-risk $\{k^{\Pi}_{t}\}:=\{\frac{{\pi}^{+}_{t}-\Pi^{+}_{t}}{\sigma^{\pi}_{t}}\}$ (see (19) or (11)) satisfies:
$k^{\Pi}_{t}=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t},\text{ }\forall t<\infty,$ (23)
with $\sigma^{\Pi}_{t}:=({\Pi}^{+}_{t}\underline{\Pi}^{+}_{t})^{\frac{1}{2}}$ and $K:=\frac{O_{f}[{\pi}^{+}_{0}]}{O_{f}[{\Pi}^{+}_{0}]}=\frac{O_{f}[{\pi}^{+}_{t}]}{O_{f}[{\Pi}^{+}_{t}]}$, a conserved quantity.
Remark 7.
This follows from Bayes’ Rule (3). Price-of-model-risk $k^{\Pi}_{t}$ is positive iff. $K>1$. It takes value $k_{\frac{1}{2}}:=(K-1)/(K+1)$ at peak risk $\sigma^{\pi}_{t}=\frac{1}{2}$, where model-risk pricing offers a risk-premium of $\frac{1}{2}k_{\frac{1}{2}}{S}^{\Delta}(t)$ (see (18)) and a gain-to-loss ratio [^5] Gain-to-loss with respect to $B$-outcomes is always $O_{f}[{\Pi}^{+}_{t}]^{-1}$, but $O_{f}[{\pi}^{b}_{t}]=1$ at peak $B$-risk, hence the result. of $(1+k_{\frac{1}{2}})/(1-k_{\frac{1}{2}})\equiv K$. To any buyers (sellers), the higher (lower) $K$ is, the better; it is natural in theory and practice to set $K\in(1,2)$ for ’competitively priced’ model-risk, with $K-1\not\gg 1$ in any case.
Remark 8.
The difference process $\{{\pi}^{+}_{t}\}-\{{\Pi}^{+}_{t}\}=\{k^{\Pi}_{t}\sigma^{\pi}_{t}\}$ has a fixed sign. Unless the model-risk is consistent (see Item-2, Section 1.1, and (14)), viable model-risk premia $\{B\_RP_{t}\}$ (see (18)) must be negative sometimes and so correspond to no known economic risk-pricing theories.
2.2.2 Connection with Familiar Price-Dynamics
For better focus, let any $B$-sure risk-premia be $B$-independent henceforth, to be written as $\check{rp}(t)=\int_{t}^{T}\check{R}_{s}ds>0$, with $\check{R}_{t}$ denoting the $B$-sure risk-premium drift at any $t<T$; note then $Y^{\Delta}(t)=S^{\Delta}(t)=\int_{t}^{T}r^{Z\Delta}_{s}ds$ as a result. Canonical pricing (22) becomes:
$S_{t}=S^{(0|\Pi)}_{t}-\check{rp}(t)=Y^{-}_{t}+S^{\Delta}(t)\Pi^{+}_{t}-\check{rp}(t),$ (24)
$dS_{t}(B)=dY^{-}_{t}(B)+S^{\Delta}(t)d\Pi^{+}_{t}(B)-r^{Z\Delta}_{t}\Pi^{+}_{j}(B)dt+\check{R}_{t}dt.$ (25)
By setup (Section 1.1) and inference dynamics (A.2-A.3), process $\{d\Pi^{+}_{t}\}$ has two sources of noise, one from $\{Z_{t}\}$, given by $\{\sigma^{lZ}_{t}dw^{Z}_{t}\}$, with $\sigma^{lZ}_{t}:=r^{Z\Delta}_{t}/\sigma^{Z}_{t}\ll 1$ (’signal-to-noise’), one from independent $B$-informative data $\{D_{t}\}$, modelled in parallel ((2)), which can be absorbed through $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$ and $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$; then:
$d\Pi^{+}_{t}(B)=\mu^{\Pi,B}_{t}dt+\sigma^{\Pi,l}_{t}dw_{t},\text{ with},$ (26)
$\mu^{\Pi,B}_{t}:=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot\big{(}{{\mathbf{1}_{+}}}(B)-{\Pi}^{+}_{t}(B)\big{)},$ (27)
$\sigma^{\Pi,l}_{t}:=(\sigma^{\Pi}_{t})^{2}\sigma^{l}_{t}.$ (28)
The above has a drift by design relative to reference belief $\{{\pi}^{+}_{t}\}$ (with dynamics (A.3)):
$\mu^{\Pi,\pi}_{t}:=\Sigma_{B}{\pi}^{B}_{t}{\mu^{\Pi,B}_{t}}=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot(\pi^{+}_{t}-\Pi^{+}_{t})=\sigma^{\Pi,l}_{t}\sigma^{l}_{t}\cdot k^{\Pi}_{t}\sigma^{\pi}_{t}\geq 0;$ (29)
note ${\mu^{\pi,\pi}_{t}}\equiv 0\equiv{\mu^{\Pi,\Pi}_{t}}$, reflecting the martingale property of inference and viable pricing. Further, as inference volatility $\sigma^{\Pi,l}_{t}$ (28) drives price-volatility (see (25-26)), price-of-model-risk $\{k^{\Pi}_{t}\}$ (by (23)) peaks where price-volatility peaks, ceteris paribus.
Pulling it together, the full asset-price dynamic, a $(\Pi^{B}_{0}\mathbf{W}^{T}_{B})$-martingale plus drift $\{\check{R}_{t}\}$, reads:
$dS_{t}(B)=$ $\big{[}\check{R}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}$ (30)
$+$ $\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{j}(B)\big{)}\cdot r^{Z\Delta}_{t}dt\big{]}$ (31)
$+$ $S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}\big{(}\mathbf{1}_{+}(B)-\Pi^{+}_{t}(B)\big{)}\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}.$ (32)
Remark 9.
Practical cases often see $B$-sure dynamics (30) and model-drift (31) dominated by model-inference (32): $\check{R}_{t}\ll S^{\Delta}(t)$ and $r^{Z\Delta}_{t}\ll S^{\Delta}(t)$; indeed firmed-up values $\{Z_{t}\}$ are often nil or fixed so that asset-pricing is just a matter of model-risk inference, with $\{dS_{t}\}$ reduced to the inferential test equation of Peskir and Shirayev (2006) (for constant $\{r^{D\Delta}_{t}\}$ and $\{\sigma^{D}_{t}\}$). Further, it often holds that relative variance $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}\ll 1$ and so $(\sigma^{l}_{t})^{2}\approx(\sigma^{lD}_{t})^{2}$: model-risk is the prime price-mover, not tangible value. Dispersion among such assets is thus driven mainly by their differing underlying model-risk inference processes.
Averaged under model-risk belief $\{{\pi}^{+}_{t}\}$, dominant price-dynamic (32) has this familiar form:
$\sigma_{t}\bigg{(}\frac{\mu_{t}}{\sigma_{t}}dt+dw_{t}\bigg{)},$ (33)
with volatility $\sigma_{t}:=\sigma^{\Pi,l}_{t}S^{\Delta}(t)$, drift $\mu_{t}:=\mu^{\Pi,\pi}_{t}S^{\Delta}(t)$, and price-of-diffusion-risk $\mu_{t}/\sigma_{t}$ ((23)):
$\frac{\mu_{t}}{\sigma_{t}}=(K^{\frac{1}{2}}-K^{-\frac{1}{2}})\sigma^{\Pi}_{t}\sigma^{\pi}_{t}\sigma^{l}_{t}=\frac{K^{\frac{1}{2}}-K^{-\frac{1}{2}}}{S^{\Delta}(t)}\bigg{(}\frac{\sigma^{\pi}_{t}}{\sigma^{\Pi}_{t}}\bigg{)}\sigma_{t}.$ (34)
Expansion in (K1)(K-1), given Remark 7, recovers the standard and ubiquitous price-of-risk vs price-volatility relationship:
$\frac{\mu_{t}}{\sigma_{t}}\approx\bigg{(}\frac{K-1}{S^{\Delta}(t)}\bigg{)}\sigma_{t}\text{ and }\mu_{t}\approx\bigg{(}\frac{K-1}{S^{\Delta}(t)}\bigg{)}\sigma_{t}^{2}.$ (35)
Remark 10.
The proportionality factor characterises ex-ante pricing through an intuitive and direct metric of asset-pricing (Remark 7). How its level comes about is a matter of economic theory (e.g. risk-aversion). Regardless, our derivation links it firmly to ex-ante beliefs, making the above relevant to bias studies and measurement under model-uncertainty.
3 Bias and Risk-Pricing Measurement via Price Anomalies
3.1 The Joint-Hypothesis Issue
To focus on main issues, as in Section 2.2.2, consider usual-case assets under model-risk $B$ with $B$-independent $B$-sure risk-premia $\{\check{rp}(t)\}$. Let $\{\mathbf{p}^{B}_{t}\}$ be the $\{\mathbf{F}_{t}\}$-adapted true conditional probabilities of model-risk outcomes, given the true and known $B$-sure laws $\mathbf{W}^{T}_{B}$ and the true but potentially unknown unconditional law $\mathbf{p}^{B}_{0}$ of model-risk $B$. Its ex-post price-drift to be realised on average $\{rp_{t}\}:=\{\Sigma_{B}\mathbf{p}^{B}_{t}{Y^{B}_{t}}\}-\{S_{t}\}$ satisfies:
$rp_{t}-\check{rp}(t)=B\_RP_{t}+(\mathbf{p}^{+}_{t}-{\pi}^{+}_{t}){S}^{\Delta}(t),\text{ }t<T.$ (36)
The LHS terms are each observable or known, but the RHS are joint: no study of risk-pricing $\{B\_RP_{t}\}$ can be done without one on bias $\{\mathbf{p}^{b}_{t}\}-\{{\pi}^{b}_{t}\}$.
The above, in standard model-risk study settings, with the ’asset’ set to ’equity index’ and ’bias’ to nil, expresses the usual attribution of ’excess historical equity-return’ (LHS) to ’model-risk premium’ (RHS). Our focus is bottom-up and when bias is not necessarily nil.
3.2 Bias and Risk-Pricing Parameters under Model-Uncertainty
An usual-case asset, if model-sure, is classical (a drift with a static noise). Upon the triggering of model-risk, it may enter a new state (e.g. from $B=+$ to $B=$-) with true probability $\mathbf{p}^{(\cdot)}_{0}$: denote the new state by ’1’, status quo, ’0’, and status quo model-sure pricing, $\{S^{\text{{\bf{0}}}}_{t}\}$.
The triggering events may or may not follow known laws. The process of how/when they arise is ignored, assuming that the average time between model crises, say some $\lambda^{-1}\gg 1$, is so long that each can be considered one-off. The period of interest is right after the arrival of a trigger but long before the next, so $t\ll\lambda^{-1}$, and long before horizon/maturity, so $t\ll T$.
Canonical asset-pricing (24), with the direction $sign[1{\text{{\bf{0}}}}]:=sign[Y^{1}_{t}-Y^{\bf{0}}_{t}]$ of potential change explicit and the $t$-dependence of impact $S^{\Delta}(t)$ dropped given $t\ll T$ (and so $S^{\Delta}(t)\gg r^{Z\Delta}_{t}$), reads:
$S_{t}=S^{\text{{\bf{0}}}}_{t}+sign[1{\text{{\bf{0}}}}]S^{\Delta}\cdot\Pi^{1}_{t},$ (37)
where the RNE belief process $\{\Pi^{1}_{t}\}\sim\{\pi^{1}_{t}\}$ is characterised by parameter $K:=(\frac{O_{f}[\pi^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]})^{sign[1{\text{{\bf{0}}}}]}$ (Corollary 1); let the risk-pricing be competitive, so $K\in(1,2)$ (Remark 7).
The relationship between the true conditional probabilities $\{\mathbf{p}^{1}_{t}\}$ (given the true unconditional law $\mathbf{p}^{1}_{0}$), reference beliefs $\{\pi^{1}_{t}\}$ and RNE beliefs $\{\Pi^{1}_{t}\}$, all $\{\mathbf{F}_{t}\}$-adapted, can be characterised by:
$\rho:=\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\pi}^{1}_{0}]},\text{ and}$ (38)
$\rho K^{sign[1{\text{{\bf{0}}}}]}\equiv\frac{O_{f}[\mathbf{p}^{1}_{0}]}{O_{f}[{\Pi}^{1}_{0}]}.$ (39)
If the change-risk is diversifiable, so unpriced, then $K=1$; if it is classical, so its law $\mathbf{p}^{1}_{0}$ is known, then $\rho=1$. Parameter $\rho$ stems from model-uncertainty and is thus unknown ex-ante.
The joint-hypothesis problem takes a concrete form: observed prices reflect product (39), so external criteria seem necessary to separate ’intention’ $K$ from ’mistake’ $\rho$. We show that this need not be so. The case of interest for us is when change is unlikely, $\mathbf{p}^{1}_{0}<\frac{1}{2}$, but is priced as rare so that (39-38) are both sizeable ($\rho\gg 1$). Such a strong status quo forms a clear base against which ’change’ is defined and assessed. Unjustified strong status quo (i.e. $\mathbf{p}^{1}_{0}\geq\frac{1}{2}$ and $\rho\gg 1$) would allow easy wins for ’bet on change’ and so should not persist.
3.3 An Ideal Model-Uncertainty Driven Market
Consider a model-uncertainty driven market, of a large number of assets, each facing some change-risk $B$, whose nature may vary from asset to asset (e.g. new CEO, tort, war) but whose true $B$-sure laws are known. Let the $B$-risks be priced (i.e. correlated with the market’s stochastic-discount-factor), for their effects on aggregate value/volatility/cross-sectional dispersion [^6] Dispersion within members of an equity index, as a priced factor, has been linked to market uncertainty, the business cycle and structural shifts (e.g. Demirer and Jategaonkar (2013), Kolari et al. (2021)).. Further, let them be economically priced, so that by Proposition 1 their pricing is canonical with respect to some reference inferential belief process of the relevant $B$-risk.
Finally, let the realisation and detection of individual change-outcomes be stochastic and independent (even for macroeconomic triggers such as war or recession). Note that such a market, when seen as a whole, is an asset with non-binary model-uncertainty; this is so even when there is only one binary macroeconomic model-risk trigger.
For brevity, let the unconditional all-cause probability of positive/negative economic change be symmetrical: at any time half of all potential changes, of any origin, are positive/negative (unconditional asymmetry is addressed in Remark 14&15). Let ’pricing intention’ $K$ ((39)) and ’apparent bias’ $\rho$ ((38)) be indifferent to the economic direction and origin of change-risk, and the size of $K$, $\rho$ and risk-impact $S^{\Delta}$, uniform across the assets. Such uniformity in practice amounts to asset dispersions being driven by differences in the paths of model-risk inference rather than in any other aspect (Remark 9). Similarly, let all risk-inference be based on log-LR processes that are homogenous in time and uniform across.
3.4 Conditional Cross-Sectional Trading
Pair-trading sorts assets into cohorts by cross-sectional statistics for long vs short positions. It means in our setup conditioning by $B$-inferential events $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, the driver of such statistics. Under (37-39), given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change, the $\{\Pi^{1}_{t}=v\}$-conditioned ex-post mean-excess $rp(\pm,v):=\pm(\mathbf{p}^{1}_{t}(v)-v)S^{\Delta}$ unpacks (by Bayes’ Rule (3)) to:
$rp(\pm,v)=\pm(\sigma^{v})^{2}\frac{1-(\rho K^{\pm 1})^{-1}}{v+\underline{v}(\rho K^{\pm 1})^{-1}}S^{\Delta},$ (40)
where $(\sigma^{v})^{2}:=v\underline{v}$, essentially making price-volatility (see (25-28)) the driver of excess. Note that the above is invariant under label-switching $1\leftrightarrow{\text{{\bf{0}}}}$ (i.e. $v\leftrightarrow\underline{v}$, $\leftrightarrow-$ and $\rho\leftrightarrow\rho^{-1}$).
Remark 11.
With no uncertainty, $\rho=1$, the above is by (23) the ex-ante risk-premium given $\{\Pi^{1}_{t}=v\}$: in classical markets all such strategising merely earns what the market deems fair.
Conditioning, or in practice, ranking, by events $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, amounts to, under strong status quo, ranking by momentum (past performance): $\{\Pi^{1}_{t}=v\}$ means $\{\Pi^{1}_{t}\}$ going from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v$ and so an excess-to-date of $\pm(v-\Pi^{1}_{0})S^{\Delta}$ ((37)). The associated model-uncertainty ($\propto(\sigma^{v})^{2}$) is a priori low (’strong status quo’) and evolves with dataflow. For efficient pair-trading, the window of opportunity is given by (C.8), Appendix C.2.
Historical drifts are hard to measure in practice (e.g. Merton (1980)), so volatility trading can be an alternative. Volatility events $\{\sigma^{\Pi}_{t}=\cdot\}$ have the form $\{\Pi^{1}_{t}=v\}\cup\{\Pi^{1}_{t}=\underline{v}\}$, $v\in(0,\frac{1}{2}]$, by (25-28). Such $\{\sigma^{\Pi}_{t}=v\underline{v}\}$-conditioned ex-post mean-excess $rp_{t}^{\sigma}(v)$ is an average over the $sign[1{\text{{\bf{0}}}}]$ of potential change and over the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-events in the volatility cohort:
$rp_{t}^{\sigma}(v):=\frac{\sum_{\pm}\mathbb{P}_{t}(\pm,v)rp(\pm,v)+\sum_{\pm}\mathbb{P}_{t}(\pm,\underline{v})rp(\pm,\underline{v})}{\mathbb{P}_{t}(v)+\mathbb{P}_{t}(\underline{v})},$ (41)
where $\mathbb{P}_{t}(\cdot)=\sum_{\pm}\mathbb{P}_{t}(\pm,\cdot):=\sum_{sign[1{\text{{\bf{0}}}}]}\mathbb{P}_{t}\big{(}\{sign[1{\text{{\bf{0}}}}]=\cdot\}\cap\{\Pi^{1}_{t}=\cdot\}\big{)}$.
The relevant likelihoods (ratios) are determined by the underlying log-LR dynamics, whose cross-sections are Normal, making closed expressions possible. As a by-product, these probability ratios predict an association of volatility with ’negative momentum’ (Remark 14), a well-known phenomenon (e.g. Ang et al. (2006) and Wang and Xu (2015)).
Observe that both the risk-reward of momentum (40) and volatility trading (41) are driven by $(\sigma^{v})^{2}$ essentially, and are thus concave as a function of $v$ and vanishing as $(\sigma^{v})^{2}\to 0$.
3.4.1 Momentum Effect and Status Quo Bias
Conditional excess (40) exhibits Momentum, the well-known anomaly (Jegadeesh and Titman (1993)): for $\rho\gg 1$, any excess-to-date $\pm(v-\Pi^{1}_{0})S^{\Delta}$, due to RNE beliefs $\{\Pi^{1}_{t}\}$ rising from $\Pi^{1}_{0}\ll\frac{1}{2}$ to $v>\Pi^{1}_{0}$, persists on average in the same direction. The effect grows with $\rho$, which by its definition (38) is an indicator of bias against change.
Momentum applies also to $\rho\ll 1$: past excess due to RNE beliefs $\{\Pi^{1}_{t}\}$ falling from $\Pi^{1}_{0}\approx 1$ to $v<\Pi^{1}_{0}$ means an average future excess of the same sign; indeed, label-switching defines a new parameter $\rho^{-1}>1$. Momentum thus vanishes only when $\rho=1$ (e.g. when there is no model-uncertainty).
Momentum trading (40) earns its best realisable excess $rp^{\pm mo}_{max}$ at $v^{\pm mo}_{max}$:
$v^{\pm mo}_{max}:=\frac{1}{(\rho K^{\pm 1})^{\frac{1}{2}}+1},\text{ }rp^{\pm mo}_{max}:=\pm\frac{(\rho K^{\pm 1})^{\frac{1}{2}}-1}{(\rho K^{\pm 1})^{\frac{1}{2}}+1}S^{\Delta}.$ (42)
Peak-profitability reads $\frac{1}{2}(rp^{+mo}_{max}-rp^{-mo}_{max})$, by pairing cohorts of conditioning-momenta $v^{\pm mo}_{max}$. As a function of bias $\rho$, risk-reward (40) and its peak (42) imply rising rewards from rising momentum up to a point, beyond which profitability falls, as observed in Ang et al. (2006).
3.4.2 Low-Risk Effect and Status Quo Bias
Low-Risk Effect (see e.g. Ang et al. (2006)) turns the classical dictum ’high risk high reward’ on its head. Momentum risk-reward (40) already hints at it: peak-reward (42) is in the low-volatility zone for large bias $\rho\gg 1$. The effect is stronger under volatility-conditioning (41), whose peak-location and -size can be found by usual methods (Appendix C). The solution at $\rho=1$ (no bias) and the leading-order solutions at $\rho\gg 1$ (high bias) are revealing and happen to have the same expression; the peak-location $v^{\sigma}_{max}$ and -size $rp^{\sigma}_{max}$ are given by:
$v^{\sigma}_{max}:=\frac{1}{\rho+1},\text{ }rp^{\sigma}_{max}:=\frac{1}{2}\frac{K-1}{K+1}S^{\Delta}.$ (43)
At $\rho=1$ the above is exact, by which $v^{\sigma}_{max}=\frac{1}{2}$, where price-volatility peaks, thus confirming ’high risk high reward’ in the absence of model-uncertainty or bias. There is a low-risk effect otherwise: peak-reward occurs at $v^{\sigma}_{max}\approx\rho^{-1}$, while peak volatility ($v=\frac{1}{2}$) is poorly rewarded ex-post: $rp^{\sigma}(\frac{1}{2})=\frac{S^{\Delta}}{2}\mathcal{O}(\frac{K-1}{\rho}-1)$.
Note the separation: bias $\rho$ is revealed by peak-location $v^{\sigma}_{max}$, and risk-pricing $K$, by peak-reward $rp^{\sigma}_{max}$, which is exactly that set ex-ante for peak-risk (Remark 7). No external speculation is needed to tell ’intended pricing’ from ’unintended bias’.
4 Summary and Discussion
We study the risk-neutral equivalent formulation of model-risk pricing. Explicit price dynamics under binary model-risks with potentially unknown laws are derived.
The results show that model-risk can be a main source of excess value, with a key role in Status Quo Bias and anomalies such as Momentum and Low-Risk. They offer an integrated and practicable approach to these highly relevant and interconnected topics.
We arrive at the above by making use of the informational constraint imposed by ’usual’ economic asset-pricing (Remark 1 and Definition 1), to show that viable asset-pricing under model-risk may take just one canonical form (Lemma 1 and Proposition 1), which leads to a unique and intuitive model-risk pricing formula with convenient properties (Corollary 1).
The sole risk-pricing parameter is associated with a conserved quantity of inferential dynamics and is a product of ’intention’ (preference) and ’mistake’ (difference between objective and subjective laws). Ex-ante (’intended’) pricing, classical or with ambiguity-aversion/robust-control, can be captured by our parameter $K\in[1,\infty)$, (39), with $K=1$ corresponding to ’no model-risk discounts’. Its derivation reveals that it sets the intended gain-to-loss ratio of the asset at peak model-risk, suggesting $K\in[1,2)$ in competitive markets (Remark 7).
The difference between true model-risk law and the reference model-risk belief on which asset-pricing is based can be captured by our parameter $\rho\in(0,\infty)$, (38), with $\rho=1$ corresponding to ’no difference’ (the classical setting), and $\rho\not=1$, ’bias’. It gives expression to Status Quo Bias and allows its effects to be disentangled from risk-pricing (Section 3.4.1-3.4.2).
The above is interesting also in light of the inertia axiom of Bewley (2002), which says, in effect, that to update an existing decision by incorporating a new state subject to Knightian uncertainty, the status quo is kept unless an alternative that is better in all scenarios exists. For our model-risk, it suggests inertial pricing, one that is piece-wise free of model-risk pricing, jumping between alternative models when appropriate, despite data that can resolve model-risk via regular inference. Asset-prices under max-min utility already have this characteristic, albeit with a built-in pessimism absent in inertial pricing. Both are viable in our setup, as ’models’ are represented by probability laws that are equivalent before any risk-resolution (that is predictable). The theoretical and practical aspects of this topic will be examined in a follow-up study.
APPENDIX
A Properties of Inferential Hypothesis Testing
  1. 1.
    Standard Continuous-Time Inference. Given the setup and Item-5 of Section 1.1, the log-LR process of time-homogeneous Wiener data processes has the form below (Peskir and Shirayev (2006)), in terms of some constant $\sigma^{l}>0$ characterising the signal-to-noise ratio of the given data,
    $dl^{+/-}_{\tau}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l})^{2}}{2}d\tau+\sigma^{l}dw_{\tau},$ (A.1)
    with $\{w_{\tau}\}$ denoting a standard Wiener process. Any time-variation may be introduced via an absolutely continuous clock-change $t\mapsto\tau(t)$. The image process $\{l^{+/-}_{t}\}:=\{l^{+/-}_{\tau(t)}\}$ is again Wiener, but with continuous time-varying parameters (e.g. Kallsen (2006)). It and its resulting inferential belief process under Bayes’ Rule (3) read: in terms of some standard Wiener process $\{w_{t}\}$,
    $dl^{+/-}_{t}(B)=(-1)^{\mathbf{1}_{-}(B)}\frac{(\sigma^{l}_{t})^{2}}{2}+\sigma^{l}_{t}dw_{t};$ (A.2)
    $\frac{d\pi^{+}_{t}(B)}{(\sigma^{\pi}_{t})^{2}}=\big{(}\mathbf{1}_{+}(B)-\pi^{+}_{t}(B)\big{)}(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}.$ (A.3)
  2. 2.
    Regular Inference. Its defining feature is the predictable resolution of risky outcomes. In standard situations, with i.i.d data, risk-resolution occurs only at infinity. Predictable finite-time resolution can be modelled via a clock-change that brings resolution forward to some $T_{\mathcal{B}}\leq\infty$. The resulting log-LR process diverges as $t\to T_{\mathcal{B}}$, but the associated beliefs and so asset-pricing remain finite and continuous at $T_{\mathcal{B}}$ under Bayes’ Rule (3).
  3. 3.
    Non-Resolution and Novikov’s Condition. For any log-LR process based on equivalent Wiener measures on the Wiener space $\mathcal{S}$ of total data, written as $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ in the financial context, one being some RNE version of the other, the Radon-Nikodym Theorem ensures $\lim_{t\to\infty}\log\frac{d\hat{\mathbf{W}}_{B}|_{t}(\cdot)}{d\mathbf{W}_{B}|_{t}(\cdot)}<\infty$ almost surely and hence a finite total variance for the associated log-LR process of form (A.2). The resolution of $B$-outcomes (Item-5, Section 1.1) thus equates to the associated LR process (the exponentiation of (A.2)), or equivalently, (A.3), violating Novikov’s Condition, and vice versa.
  4. 4.
    Adjacency. Given any pair of log-LR process $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ of respective defining laws $\{\mathbf{W}_{+},\mathbf{W}_{-}\}$ and $\{\hat{\mathbf{W}}_{+},\hat{\mathbf{W}}_{-}\}$ such that equivalence $\hat{\mathbf{W}}_{B}\sim\mathbf{W}_{B}$ holds on the space $\mathcal{S}$ of total data, then they almost surely can differ at most by a constant as $t\to\infty$, since:
    $\hat{l}^{+/-}_{t}-l^{+/-}_{t}=\log{\frac{d\hat{\mathbf{W}}_{+}|_{t}}{d\mathbf{W}_{+}|_{t}}}-\log{\frac{d\hat{\mathbf{W}}_{-}|_{t}}{d\mathbf{W}_{-|_{t}}}},\text{ }\forall t\in\mathbb{R}^{+},$ (A.4)
    where the RHS is convergent under equivalence; this is the case even when each of $\{l^{+/-}_{t}\}$ and $\{\hat{l}^{+/-}_{t}\}$ is resolving (i.e. divergent in total variance and so total drift).
    Lemma 1.
    No two resolving log-LR processes, unless identical up to a constant difference, can be deterministically and time-homogenously related to one another.
[…]
    $g^{\prime\prime}=-(-1)^{\mathbf{1}_{+}(B)}(g^{\prime}-1)g^{\prime}.$ (A.5)
    If $B=+$, then $g^{\prime}e^{-g}=g^{\prime}(0)e^{-g(0)}e^{-\mathbf{Id}}$, so $g=g(0)-\log[g^{\prime}(0)e^{-\mathbf{Id}}+1-g^{\prime}(0)]$; we then have $g^{\prime}(0)=1$ as $\{\hat{l}^{+/-}_{t}\}$ and $\{{l}^{+/-}_{t}\}$ are resolving. If $B=$-, so $g=g(0)+\log[g^{\prime}(0)e^{\mathbf{Id}}+1-g^{\prime}(0)]$, then almost-sure divergence $\lim_{t\to\infty}\hat{l}^{+/-}_{t}=-\infty=\lim_{t\to\infty}{l}^{+/-}_{t}$ requires the same.
B PROOF OF PROPOSITION 1
[…]
Under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$, with $d\mathbf{W}^{T}_{-,t}=\sigma^{Z}_{t}dw^{Z}_{t}+\sigma^{D}_{t}dw^{D}_{t}$, $\{dw^{Z}_{t}\}$ and $\{dw^{Z}_{t}\}$ being independent standard Wiener processes (Item-3&5, Section 1.1), we have:
[…]
$(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)(A^{+}_{t})^{\prime}_{\pi}\bigg{[}($ $\mathbf{1}_{+}-\pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+(\sigma^{\pi}_{t})^{2}\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}\cdot(\sigma^{l}_{t})^{2}dt\text{ }+\text{ }\sigma^{l}_{t}dw_{t}\bigg{]},$ (B.2)
with model-drift $r^{Z\Delta}_{t}dt=-{d}Y^{\Delta}(t)$ (Item-2, Section 1.1) and $(\sigma^{l}_{t})^{2}:=(\sigma^{lZ}_{t})^{2}+(\sigma^{lD}_{t})^{2}$, where $\sigma^{lZ}_{t}:={r^{Z\Delta}_{t}}/{\sigma^{Z}_{t}}$ (’signal-to-noise’) stems from data $\{Z_{t}(B)\}$ (Peskir and Shirayev (2006)), with $\sigma^{lD}_{t}$ likewise from $B$-conditionally independent data $\{D_{t}(B)\}$, and so $\sigma^{l}_{t}dw_{t}:=\sigma^{lZ}_{t}dw^{Z}_{t}+\sigma^{lD}_{t}dw^{D}_{t}$.
Without risk-pricing, $\{A^{+}_{t}\}=\{\pi^{+}_{t}\}$, we have trivial viability under reference belief $\pi^{B}_{0}\mathbf{W}^{T}_{B}$:
$dS^{(0|\pi)}_{t}=\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})\cdot r^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)\big{[}(\mathbf{1}_{+}-\pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}.$ (B.3)
With viable model-risk only pricing, under some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}\sim\pi^{B}_{0}\mathbf{W}_{B}^{T}$, we have $\{S^{(0|A)}_{t}\}=\{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{S}^{B}_{t}\}$, and the following, by (8-11) and (19),
$\pi^{+}_{t}-A^{+}_{t}=k^{{A}}_{t}\sigma^{\pi}_{t}=k^{\hat{\pi}}_{t}\sigma^{\pi}_{t}+\frac{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{RP}^{B}_{t}}{{Y}^{\Delta}(t)}=(\pi^{+}_{t}-\hat{\pi}^{+}_{t})+\frac{\Sigma_{B}\hat{\pi}^{B}_{t}\hat{RP}^{B}_{t}}{{Y}^{\Delta}(t)}.$ (B.4)
Remark 12.
Reference model-risk beliefs $\{{\pi}^{+}_{t}\}$ and any of its RNE $\{\hat{\pi}^{+}_{t}\}$ are adjacent (Item-4, Appendix A), implying that at any fixed $t<T$ for $\sigma^{\pi}_{t}\ll\frac{1}{2}$, by (23) when $\{{\hat{\pi}}^{+}_{t}\}=\{\Pi^{+}_{t}\}$, we have $\mathcal{O}({\pi}^{+}_{t}-{\hat{\pi}}^{+}_{t})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and so $\mathcal{O}({k}^{\hat{\pi}}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$ for any RNE price-of-model-risk (11). In turn, by Lemma 1, the LHS and RHS of (B.4) being equal implies $\mathcal{O}({\hat{RP}^{+}_{t}})=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}({k}^{A}_{t})=\mathcal{O}(\sigma^{\pi}_{t})$, as well as $\mathcal{O}((A^{+}_{t})^{\prime}_{\pi})=\mathcal{O}((A^{+}_{t})^{\prime\prime}_{\pi})=\mathcal{O}(1)$ and $\mathcal{O}(d{\hat{RP}^{+}_{t}})=\mathcal{O}(d(\sigma^{\pi}_{t})^{2})=\mathcal{O}((\sigma^{\pi}_{t})^{4})$.
Consider first the case of trivial $\{D_{t}\}$, so that the drift of $\{dS^{(0|A)}_{t}\}$ is nil in expectation under some RNE measure $\hat{\pi}^{B}_{0}\hat{W}_{B}^{Z}|_{T}$ satisfying $d\hat{W}^{Z}_{B,t}=\hat{r}_{B,\text{ }t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$, $\hat{r}_{B,\text{ }t}dt=-d\hat{RP}^{B}_{t}$, $t<T$.
[…]
$\frac{-\nu_{t}}{r^{Z\Delta}_{t}}=(\sigma^{\pi}_{t})^{2}Y^{\Delta}(t)(A^{+}_{t})^{\prime}_{\pi}\frac{(\sigma^{lZ}_{t})^{2}}{r^{Z\Delta}_{t}}\bigg{[}(A^{+}_{t}-{\pi}^{+}_{t})+(\sigma^{\pi}_{t})^{2}\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}-\frac{-\nu_{t}}{r^{Z\Delta}_{t}}\bigg{]}.$ (B.5)
By Remark 12, we have $\mathcal{O}(LHS)=\mathcal{O}((\sigma^{\pi}_{t})^{2})$ and $\mathcal{O}(RHS)=\mathcal{O}((\sigma^{\pi}_{t})^{4})$, $\forall\sigma^{\pi}_{t}\ll\frac{1}{2}$, so nil-drift is possible only if each side vanishes on its own, meaning:
$\frac{(A^{+}_{t})^{\prime\prime}_{\pi}}{2(A^{+}_{t})^{\prime}_{\pi}}=\frac{{\pi}^{+}_{t}-A^{+}_{t}}{(\sigma^{\pi}_{t})^{2}}.$ (B.6)
This has solutions of the form $\{A^{+}_{t}\}=\{\Pi^{+}_{t}\}$ only, where $\{\Pi^{+}_{t}\}$ is any model-risk inference based on any measure $\Pi^{B}_{0}W_{B}^{Z}|_{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under the reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d\hat{\mathbf{W}}^{T}_{B,t}=\hat{r}_{B,\text{ }t}dt+d\mathbf{W}^{T}_{-,t}$, $\hat{{r}}_{B,\text{ }t}dt=-d\hat{{RP}}^{B}_{t}$. To conclude, the RNE measure $\{S^{(0|A)}_{t}\}$ must have the canonical form: $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Part II. The viability of canonical price process $\{S^{(RP|{\Pi})}_{t}\}=\{{S}^{(0|{\Pi})}_{t}\}-\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$.
Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}RP^{B}(t)$ that stems from $B$-sure risk-pricing:
$dS^{(RP|\Pi)}_{t}=\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\cdot\check{r}^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}+R^{(\cdot)}_{t}dt,$ (B.7)
where $\check{r}^{Z\Delta}_{t}dt:=-{d}{}S^{\Delta}(t)$, with $S^{\Delta}(t)\equiv Y^{\Delta}(t)-RP^{+}(t)+RP^{-}(t)$ (by (12-14). As such, it is viable if drift $\{R^{B}_{t}dt\}$ can be ’absorbed’ into the known $(\Pi^{B}_{0}\mathbf{W}_{B}^{T})$-martingale $\{d{S}^{(0|{\Pi})}_{t}\}$, that is, whenever
$\frac{{R^{\pm}_{t}}/{S^{\Delta}(t)}}{{\sigma^{Z}_{t}}/{S^{\Delta}(t)}+{(\sigma^{\Pi}_{t})^{2}{\sigma^{lD}_{t}}\sqrt{1+\text{F}^{Z/D}_{t}}}}$
is square-integrable up to horizon $T$ (Novikov’s Condition), where $(\sigma^{l}_{t})^{2}\equiv(\sigma^{lD}_{t})^{2}(1+\text{F}^{Z/D}_{t})$ is written in terms of the relative variance $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}$ of data $\{Z_{t}\}$ to $\{D_{t}\}$.
C THE MOMENTUM AND LOW-RISK FORMULAE
For the assets of Section 3.2-3.3, all the distributions required stem from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
$\mathcal{N}\bigg{(}\mu^{l}_{[t]}(B)=\frac{(-1)^{B+1}}{2}(\sigma^{l})^{2}t,\text{ }(\sigma^{l}_{[t]})^{2}=(\sigma^{l})^{2}t\bigg{)},\text{ }B\in\{0,1\}.$ (C.1)
The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices (37) then derive via (3) by change of variable: for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, we have $dl=(v\underline{v})^{-1}dv$, with:
$l=H_{\Pi\pm}+\log\circ\text{ }O_{f}[v],$ (C.2)
$H_{\Pi\pm}:=\log\circ\text{ }O_{f}[\underline{\Pi}^{1}_{0\pm}]=H_{\mathbf{p}}+\log(\rho K^{\pm 1}),$ (C.3)
$H_{\mathbf{p}}:=\log\circ\text{ }O_{f}[\underline{\mathbf{p}}^{1}_{0}],$ (C.4)
where recall (39), $O_{f}[\Pi^{1}_{0\pm}]=O_{f}[\mathbf{p}^{1}_{0}]/(\rho K^{\pm 1})$ given $sign[1{\text{{\bf{0}}}}]=\pm$. The $H$-variables above are inferential milestones: $H_{\Pi\pm}$ is the log-LR hurdle for event $\{\Pi^{1}_{t}\geq\frac{1}{2}\}$, and $H_{\mathbf{p}}$, for $\{\mathbf{p}^{1}_{t}\geq\frac{1}{2}\}$.
C.1 Tracking Inferential Progress and Bias Dominance
The degree of $B$-certainty as data accumulate may be assessed in the standard way for Normal distributions. Its objective level is tracked by some $|C^{\mathbf{p}}_{t}|$, where, with $t_{\mathbf{p}}:=2H_{\mathbf{p}}/(\sigma^{l})^{2}$,
$C^{\mathbf{p}}_{t}(B):=\frac{H_{\mathbf{p}}-\mu^{l}_{[t]}(B)}{\sigma^{l}_{[t]}}=\frac{1}{2}\sigma^{l}_{[t]}\cdot\big{(}{(-1)^{B}}+\frac{t_{\mathbf{p}}}{t}\big{)}.$ (C.5)
Certainty level $|C^{\mathbf{p}}_{t}|$ is high if $\sigma^{l}_{[t]}\ll 1$ (little data) or $\sigma^{l}_{[t]}\gg 1$ (lots of data); it bottoms out at 0 if and when $H_{\mathbf{p}}=\mu^{l}_{[t]}(B)$ has a solution, that is, if and when $t=(-1)^{B+1}t_{\mathbf{p}}$ and $B=1$.
For the price-implied RNE inference process $\{\Pi^{1}_{t}\}$, the same indicator, denoted $|C^{\Pi}_{t\pm}|$, reads:
$C^{\Pi}_{t\pm}(B):=$ $\frac{H_{\Pi\pm}-\mu^{l}_{[t]}(B)}{\sigma^{l}_{[t]}}=\frac{1}{2}\sigma^{l}_{[t]}\cdot\big{(}{(-1)^{B}}+\frac{t_{\Pi\pm}}{t}\big{)}$ (C.6)
$=$ $\frac{1}{2}\sigma^{l}_{[t]}\cdot(\frac{t_{\rho}}{t}\pm\frac{t_{K}}{t})+C^{\mathbf{p}}_{t}(B),$ (C.7)
where $t_{\Pi\pm}=t_{\rho}\pm t_{K}+t_{\mathbf{p}}$, with $t_{\rho}:={2\log(\rho)}/{(\sigma^{l})^{2}}>0$ and $t_{K}:={2\log K}/{(\sigma^{l})^{2}}>0$ (recall $K-1\not\gg 1$ in general and $K\in(1,2)$ in competitive markets). The pattern of $|C^{\Pi}_{t\pm}|$ parallels that of $|C^{\mathbf{p}}_{t}|$; it has 0 as a minimum if and when $t=(-1)^{B+1}t_{\Pi\pm}>0$.
Bias creates a ’burden of proof’ $t_{\rho}$, and $t_{\rho}/t$ tracks how it is overcome. We consider cases of positive objective hurdle $H_{\mathbf{p}}>0$ and bias domination $H_{\Pi\pm}-H_{\mathbf{p}}\gg 0$, so $t_{\mathbf{p}}>0$ and $t_{\rho}\gg t_{K}$, where ’objective burden’ $t_{\mathbf{p}}$ is ’non-extreme’, to exclude easy inference or profits.
C.2 The Window of Opportunity
Excess-profit opportunities occur when data become sufficient for the conditional probabilities of change to be meaningful but insufficient for subjective inference to overcome bias:
$\big{\{}\frac{t_{\mathbf{p}}}{t}\ll 1\big{\}}\text{ }\bigcap\text{ }\big{\{}\frac{t_{\rho}}{t}\gg 1\big{\}};$ (C.8)
the first demand means data dominance over objective hurdle and the second, bias over data.
Remark 13.
Consider profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ and $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ (see (43)), with $t$ fixed inside the window above and $u>0$ large enough for the second event to be outside the window. Their respective probability densities are $\propto(t)^{-1/2}e^{-|C^{\mathbf{p}}_{t}|^{2}/2}$ and $\propto(t\pm u)^{-1/2}e^{-|C^{\mathbf{p}}_{(t\pm u)}|^{2}/2}$. Thus, in-window events dominate, with a log-LR of $\log[1\pm u/t]+|C^{\mathbf{p}}_{(t\pm u)}|^{2}-|C^{\mathbf{p}}_{t}|^{2}\gg 1$, which diverges as $(1-u/t)\to 0$ or $(1+u/t)\to\infty$, $t$ fixed.
C.3 Momentum and Volatility Mixtures
The probabilities relevant to volatility-conditioned mean-excess (41) obey the rule:
[…]
with $\mathbb{P}_{t}(\Pi^{1}_{t}=v|\text{ }\pm)$ from Normal distribution (C.1) via (C.2). The likelihood-ratio of events $\{\Pi^{1}_{t}=\underline{v}\}$ vs $\{\Pi^{1}_{t}=v\}$, $v\in(0,\frac{1}{2}]$, given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change, is:
$R_{t|\pm}(\underline{v}/v)(B):=\frac{\mathbb{P}_{t}(\underline{v}|\pm)}{\mathbb{P}_{t}(v|\pm)}(B)=({\underline{v}}/{v})^{-(t_{\Pi\pm}/{t})-(-1)^{B}},\text{ }B\in\{0,1\}.$ (C.10)
Note that $\{\Pi^{1}_{t}=v\}$-events dominate when bias does (i.e. $t_{\rho}/{t}\gg 1$), regardless of $B$-outcomes.
To compute (41) we also need the mix-ratio function $M_{t|v}:={\mathbb{P}_{t}(+|v)}/{\mathbb{P}_{t}(-|v)}$ of the event-set $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, between member-events with $\{sign[1{\text{{\bf{0}}}}]=+\}$ and with $\{sign[1{\text{{\bf{0}}}}]=-\}$. Given (C.9), we have $M_{t|v}={\mathbb{P}_{t}(v|+)}/{\mathbb{P}_{t}(v|-)}$, since $\mathbb{P}_{t}(\pm)=\frac{1}{2}$ by setup. Hence:
$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}K^{-(t_{\mathbf{p}}/t)-(-1)^{B}},$ (C.11)
and likewise the volatility-conditioned ratio $M_{t|v\underline{v}}:={\mathbb{P}_{t}(+|v\underline{v})}/{\mathbb{P}_{t}(-|v\underline{v})},\text{ }v\in(0,\frac{1}{2}]$:
$M_{t|v\underline{v}}=M_{t|v}\frac{1+R_{t|+}(\underline{v}/v)}{1+R_{t|-}(\underline{v}/v)}.$ (C.12)
Note the source of the conditionalities above, model-risk pricing $K$. For $K=1$, unpriced model-risk, conditional mixes equal the unconditional. For $K\in(1,2)$, competitively priced model-risk, conditional mixes are perturbations of the unconditional.
Remark 14.
Both mix-ratios (C.11-C.12) are monotone declining in $v$, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$. That is, for $\rho\gg 1$, in the window (C.8) of opportunity, when trading events are most abundant, the observed mix at high volatility can be highly negative vs the unconditional level. Mix reverts to its unconditional level as $t\to\infty$ when model-risk resolves. Any uneven unconditional background merely introduces a constant factor in (C.11-C.12); all above statements remain valid.
C.4 Peak-Reward Location and Size for Volatility-Conditioned Trading
Focusing first on the $\{\Pi^{1}_{t}=v\}$-contributions to volatility-conditioned mean-excess (41), given mix-function (C.11) and condition (C.8), it has a unique optimum, (43), with $\mathcal{O}(\frac{(K-1)H_{\mathbf{p}}}{(\sigma^{l})^{2}t})$-errors; the accuracy of solution (43) improves with data accumulation.
Further, given (C.10) and under condition (C.8), the $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are no more than $\mathcal{O}(\rho^{-t_{\rho}/{t}})$. That is, the leading order solution (43) is not affected by these contributions whenever $t_{\rho}/t\gg 1$, i.e. while bias dominates.
Remark 15.
For uneven unconditional-mix, $\mathbb{P}_{t}(\pm)\not=\frac{1}{2}$, the same features apply. For unevenness that tilts negatively (positively), the location of peak-excess moves left (right) vs the even solution (43), and its size, down (up).
Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are invariant under $v\leftrightarrow\underline{v}$ switching (see (40) and (C.10-C.12)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming peak-volatility as the location of peak-reward and ’high risk high reward’ under classical conditions.